### v2

# On the classical solution for the steady triple-deck equations

Abstract.
This paper establishes the existence and uniqueness of classical solutions to the steady Triple-Deck equations, which describe incompressible boundary layer flow over localized roughness at high Reynolds numbers. The triple-deck theory was developed to overcome the Goldstein singularity in classical Prandtl boundary layer theory, capturing the interaction between the viscous sublayer, main layer, and upper layer when surface roughness of height $O({\rm Re}^{-\frac{5}{8}})$ is present. The key ingredients of this paper include: (1) A decomposition separating roughness effects from nonlinear difficulties; (2) A novel Green’s function using Airy functions that overcomes low-frequency singularities via the non-vanishing property of $\sqrt{3}Ai(z)+Bi(z)$; (3) The introduction of weighted Sobolev norms ${\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big\|}_{L^{2}}$ of the vorticity yielding $M$-independent estimates for displacement $A$. As a byproduct, local uniqueness of Couette flow is established when $F=0$.
1. Introduction
1.1. The Triple-Deck equations
In 1904, Prandtl [21] proposed the boundary layer theory to address attached flow problems, achieving remarkable success in fluid dynamics research. However, when dealing with flow separation induced by adverse pressure gradients, this theory encounters singularity issues. Notably, Landau and Lifshitz [14] revealed that the velocity component normal to the body surface grows unboundedly (termed the Goldstein singularity), which hinders the solution from being further extended to the reverse-flow region [6]. To overcome these difficulties, scholars including Messiter [17] and Stewartson [27] developed the triple-deck theory from the late 1960s to the early 1970s. In this theory, three distinct layers—a viscous sublayer, a main layer, and an upper layer—appear in the transverse direction, for which uniformly valid approximate solutions are derived by the asymptotic matched method.
The 2D steady incompressible Navier-Stokes equations take the following dimensionless form:
(1.1) $\left\{\begin{aligned} &{\mathbf{u}}_{NS}\cdot\nabla{\mathbf{u}}_{NS}-\epsilon^{8}\Delta{\mathbf{u}}_{NS}+\nabla p_{NS}=0,\\ &\nabla\cdot{\mathbf{u}}_{NS}=0.\end{aligned}\right$.
Here $(u_{NS},v_{NS})$ denotes the velocity vector, $p_{NS}$ pressure, and $\epsilon={\rm Re}^{\frac{1}{8}}$ with the Reynolds number ${\rm Re}\gg 1$. In physics, the triple-deck structure emerges when localized surface roughness of height ${\mathcal{O}}(\epsilon^{5})$ is present. We then denote it as
$\Big\{(X,Y)\big|\ Y=\epsilon^{5}F\big(\frac{X-1}{\epsilon^{3}}\big),\ X\in{\mathbb{R}}\Big\}$
for some given function $F={\mathcal{O}}(1)$ as $\epsilon\to 0$. Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the ${\mathcal{O}}(\epsilon^{3})$ vicinity of the roughness (see figure 1).
Figure 1. Demonstration of the triple-deck structure (not to scale): I Viscous Sublayer; II Main Layer; III Upper Layer; $\delta$ Thickness of Prandtl layer.
We demonstrate the scales of $(u_{NS},v_{NS},p_{NS})$ in Table 1.
I II III
$u_{NS}$ ${\mathcal{O}}(\epsilon)$ $\tilde{U}_{0}(\frac{Y}{\epsilon^{4}})+{\mathcal{O}}(\epsilon)$ $1+{\mathcal{O}}(\epsilon^{2})$
$v_{NS}$ ${\mathcal{O}}(\epsilon^{3})$ ${\mathcal{O}}(\epsilon^{2})$ ${\mathcal{O}}(\epsilon^{2})$
$p_{NS}$ ${\mathcal{O}}(\epsilon^{2})$ ${\mathcal{O}}(\epsilon^{2})$ ${\mathcal{O}}(\epsilon^{2})$
Table 1. Scales of velocities and pressure in the triple-deck structure: $\tilde{U}_{0}$ Streamwise velocity in Prandtl layer.
Based on the scaling relations in Figure 1 and Table 1, and via the method of matched asymptotic expansions applied to (1.1) (see Appendix for details), we derive the governing equations for the viscous sublayer (region I) as
(1.2) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0.\end{aligned}\right$.
Here $y_{*}=\frac{Y}{\epsilon^{5}}$ and $(\epsilon u_{*},\epsilon^{3}v_{*},\epsilon^{2}P_{*})$ are the leading-order terms of $(u_{NS},v_{NS},p_{NS})$. The velocity field is subject to the no-slip, non-penetration conditions,
(1.3) $(u_{*},v_{*})\big|_{y_{*}=F(x_{*})}=0.$
In the upstream limit, the velocity approaches the unperturbed Blasius profile,
(1.4) $(u_{*}-y_{*})\big|_{x_{*}\downarrow-\infty}=0,\quad P_{*}\big|_{x_{*}\downarrow-\infty}=0.$
To close the system, the upper boundary condition of the sublayer equations is enforced by matching with the main layer solution (see Appendix for details). Then we can derive the leading-order pressure as follows
(1.5) $P_{*}=-\frac{1}{\pi}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta.$
And the upper boundary condition of $u_{*}$ is
(1.6) $(u_{*}(x_{*},y_{*})-y_{*})\big|_{y_{*}=f(x_{*})}=A_{*}(x_{*}).$
Here $f$ is given (to be determined later). Consequently, by (1.2)-(1.6), we derive the closed system for the viscous sublayer as
(1.7) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\quad F(x_{*})<y_{*}<f(x_{*}),x_{*}\in{\mathbb{R}},\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\quad F(x_{*})<y_{*}<f(x_{*}),x_{*}\in{\mathbb{R}},\\ &(u_{*},v_{*})\big|_{y_{*}=F(x_{*})}=0,\quad(u_{*}(x_{*},y_{*})-y_{*})\big|_{y_{*}=f(x_{*})}=A_{*}(x_{*}),\quad x_{*}\in{\mathbb{R}},\\ &(u_{*}-y_{*})\big|_{x_{*}\downarrow-\infty}=0,\quad P_{*}\big|_{x_{*}\downarrow-\infty}=0,\quad F(-\infty)<y_{*}<f(-\infty),\\ &P_{*}=-\frac{1}{\pi}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta.\end{aligned}\right$.
The system (1.7) is referred to as the Triple-Deck equations.
Recall that the Prandtl equations of classical boundary layers take
(1.8) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\quad\text{in }(0,L_{0})\times{\mathbb{R}}_{+},\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\quad\text{in }(0,L_{0})\times{\mathbb{R}}_{+},\\ &(u_{*},v_{*})\big|_{y_{*}=0}=0,\quad x_{*}\in{\mathbb{R}}_{+},\\ &u_{*}\big|_{x_{*}=0}=u_{0},\quad y_{*}\in{\mathbb{R}}_{+},\end{aligned}\right$.
where $u_{0}$ is prescribed. Mathematically, the Triple-Deck equations share a structural resemblance to the Prandtl equations (1.8), yet they differ fundamentally in domain and pressure determination. The Prandtl equations are defined over $(0,L_{0})\times\mathbb{R}_{+}$, with pressure $P_{*}$ dictated by external potential flow $u_{E}(x_{*})$ via the Bernoulli relation
(1.9) $P_{*}+\frac{1}{2}u_{E}^{2}=\text{Const,}$
which means that the pressure is a given function. While, the pressure of Triple-Deck model is given by (1.5) with the unknown function $A_{*}$.
The existence and uniqueness of classical solutions to the Prandtl equations were first established by Oleinik [19]; subsequent decades have yielded rich results on well-posedness, asymptotic behavior, and regularity [7, 8, 16, 23, 24, 29, 30]. Moreover, flow separation and reverse-flow phenomena have attracted considerable attention from many scholars [2, 10, 15, 25].
When $F=0$, it is easy to see that $(u_{*},v_{*},P_{*},A_{*})=(y,0,0,0)$ is a trivial solution to (1.7). In a recent work [9], Iyer and Maekawa proved the rigidity of this trivial solution under certain assumptions on the solution. However, physicists found that the triple-deck structure only appears when $F\not\equiv 0$. Since the $A_{*}$ remains unknown in this case, finding a non-trivial solution becomes a highly non-trivial problem. Numerous results are available on physical analysis and numerical simulations; see [1, 4, 12, 26, 32]. Yet, if $F\not\equiv 0$, the well-posedness of (1.7) still lacks rigorous mathematical justification. For the unsteady case, Iyer and Vicol [11] and Gérard-Varet et al. [5] established well-posedness in real-analytic spaces, whereas Dietert and Gérard-Varet [3] exhibited instability for some smooth initial data, indicating that the analytic class is optimal for well-posedness.
1.2. Main results
To simplify the system (1.7), we introduce the new coordinate $(x,y)$ and new unknowns $(u,v,P,A)$:
(1.10) $x=x_{*},\ y=y_{*}-F(x_{*}),$
$(u_{*},v_{*},P_{*},A_{*})=(y+u,v+(y+u)\frac{dF}{dx},P,A).$
Then the problem (1.7) can be reduced to
(1.11) $\left\{\begin{aligned} &y\partial_{x}u+v+\partial_{x}|\partial_{x}|A-\partial_{y}^{2}u=-u\partial_{x}u-v\partial_{y}u,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &\partial_{x}u+\partial_{y}v=0,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &(u,v)\big|_{y=0}=(0,0),\quad u\big|_{y=f(x)-F(x)}=A(x)+F(x),\quad x\in{\mathbb{R}},\\ &\lim_{x\to-\infty}u=0,\quad 0<y<f(-\infty)-F(-\infty).\end{aligned}\right$.
For convenience, we take $f(x)=M+F(x)$ for $M>0$. We introduce the vorticity $\omega=\partial_{y}u$ and
(1.12) $u(x,y)=I_{y}[\omega]:=\int_{0}^{y}\omega dy^{\prime}.$
Differentiating the first equation of (1.11) in the variable $y$, we obtain the perturbation system about $(\omega,A)$ as follows
(1.13) $\left\{\begin{aligned} &y\partial_{x}\omega-\partial_{y}^{2}\omega=-u\partial_{x}\omega-v\partial_{y}\omega,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &\partial_{y}\omega\big|_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[\omega]=A(x)+F(x),\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Here $F$ is a prescribed function. We denote $\Omega_{M}={\mathbb{R}}\times[0,M]$. It is easy to see that the Couette flow $(u_{*},v_{*},P_{*},A_{*})$ corresponds to the trivial solution $(u,v,P,A)=(0,0,0,0)$ when $F=0$.
Now we introduce the functional framework of classical solutions. For the vorticity $\omega$, we define the weighted Sobolev spaces by
$X_{\bar{\alpha}}=$ $\ \bigg\{f\big|\ {\big\|f\big\|}_{X_{\bar{\alpha}}}={\big\|f\big\|}_{L^{2}(\Omega_{M})}+{\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big\|}_{L^{2}(\Omega_{M})}+{\big\|f\big\|}_{Y_{\bar{\alpha}}}<\infty\bigg\},$
$Y_{\bar{\alpha}}=$ $\ \bigg\{f\big|\ {\big\|f\big\|}_{Y_{\bar{\alpha}}}={\big\|y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}f\big\|}_{L^{2}(\Omega_{M})}+{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big\|}_{L^{2}(\Omega_{M})}$
$\qquad\qquad\qquad\quad+{\big\||\partial_{x}|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}+{\big\||\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}f\big\|}_{L^{2}(\Omega_{M})}$
$\qquad\qquad\qquad\quad+{\big\||\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}<\infty\bigg\},$
where the parameter $\bar{\alpha}$ is nonnegative and $|\partial_{x}|^{s}:=\mathcal{F}^{-1}[|\xi|^{s}\mathcal{F}[\cdot]]$ denotes the Fourier multiplier of order $s$. For the displacement function $A$, we define
$X_{\alpha,\infty}=\bigg\{f\big|\ {\big\||\partial_{x}|^{\frac{5}{6}}f\big\|}_{H^{\frac{4}{3}+\frac{\alpha}{2}}({\mathbb{R}})}+{\big\|f\big\|}_{L^{\infty}({\mathbb{R}})}<\infty\bigg\}.$
Then the functional space for $(\omega,A)$ incorporating the upper boundary condition is defined by
${\mathcal{X}}_{\alpha,M}=\bigg\{(\omega,A)\big|$ $\ {\big\|(\omega,A)\big\|}_{{\mathcal{X}}_{\alpha,M}}={\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|A\big\|}_{X_{\alpha,\infty}}<\infty,$
$\ \ \omega(\cdot,y)\big|_{y=M}=\frac{Ai(M\partial_{x}^{\frac{1}{3}})}{Ai^{\prime}(0)}\partial_{x}^{\frac{2}{3}}|\partial_{x}|A\bigg\}.$
Here $Ai(z)$ is the Airy function satisfying $\frac{d^{2}}{dz^{2}}Ai(z)-zAi(z)=0,\ Ai\big|_{z\uparrow+\infty}=0$.
To seek the solution to the system (1.13), we introduce the following crucial decomposition $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$, where $(\omega_{0},A_{0})$ and $(\bar{\omega},\bar{A})$ solve the following system
(1.14) $\left\{\begin{aligned} &y\partial_{x}\omega_{0}-\partial_{y}^{2}\omega_{0}=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\omega_{0}\big|_{y=0}=\partial_{x}|\partial_{x}|A_{0},\quad I_{M}[\omega_{0}]=A_{0}+F,\quad x\in{\mathbb{R}},\end{aligned}\right$.
and
(1.15) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}-\partial_{y}^{2}\bar{\omega}=-(u\partial_{x}\omega+v\partial_{y}\omega)\triangleq{\bar{h}},\quad(x,y)\in\Omega_{M},\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\bar{\omega}\big|_{y=0}=\partial_{x}|\partial_{x}|\bar{A},\quad I_{M}[\bar{\omega}]=\bar{A},\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Our first main result is the solvability of the linear system (1.14).
Theorem 1.1.
Let $\alpha\in(0,\frac{7}{3}]$. The linear system (1.14) admits a solution $(\omega_{0},A_{0})\in{\mathcal{X}}_{\alpha,M}$ satisfying the estimate
(1.16) ${\big\|\omega_{0}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|A_{0}\big\|}_{X_{\alpha,\infty}}\leq$ $\ C{\big\|F\big\|}_{H^{2}({\mathbb{R}})},$
where $C>0$ is independent of $\alpha,M$.
The second main result of this paper is the existence and uniqueness of classical solution to the nonlinear system (1.13).
Theorem 1.2.
Let $\alpha\in(\frac{2}{3},\frac{7}{3}]$ and $\varepsilon\in(0,\frac{1}{6}]$. There exist a small constant $\delta_{0}=\delta_{0}(\alpha,\varepsilon)>0$ and a sufficiently large constant $M_{0}$, such that when $M\geq M_{0}$ and
${\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\delta_{0}M^{-\frac{5+3\varepsilon}{2}},$
the system (1.13) admits a solution $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}$ with
${\big\|\bar{\omega}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}\big\|}_{X_{\alpha,\infty}}\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}.$
Here $C>0$ is independent of $\alpha,\varepsilon,M$. Moreover, $(\omega,A)$ is the unique solution to the system (1.13) within ${\mathcal{X}}_{\alpha,M}$ if
${\big\|(\omega,A)\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq 2\delta_{0}M^{-\frac{5+3\varepsilon}{2}}.$
Remark 1.1.
Our theoretical results align closely with the physical phenomena. The linear component $(\omega_{0},A_{0})$ captures the dominant behavior of $(\omega,A)$. Considering the scaling relation $M={\mathcal{O}}(\text{Re}^{\frac{\alpha_{0}}{8}})$ for $\alpha_{0}\in(0,\frac{3}{4})$ (see Appendix for details), our analysis shows that the allowable roughness scale ${\mathcal{O}}(\text{Re}^{-a_{*}})$ with $a_{*}\in(0,\infty)$ falls within the physically reasonable range $0,\infty)$ (see [26]).
Remark 1.2.
If $F=0$, Iyer and Maekawa [9] derived the local rigidity of Couette flow. Specifically, they proved that an $L^{2}$ strong solution $(\omega,A)$ of (1.13) is trivial if ${\big\|(\omega,A)\big\|}_{sc}$ is sufficiently small, where ${\big\|\cdot\big\|}_{sc}$ is a scaling invariant norm. As a byproduct of Theorem 1.2, we further establish the local uniqueness of Couette flow: the system (1.13) has a locally unique solution in ${\mathcal{X}}_{\alpha,M}$, which is a trivial solution.
Remark 1.3.
We adopt $M\gg 1$ in our model for consistency with physical observations. Indeed, for general $M>0$, the results may also be got by Remark 3.2.
1.3. Main difficulties and Key ideas
The system (1.13) is a degenerate nonlinear parabolic equation with variable coefficients and its degeneration occurs on the wall $y=0$. To solve the system (1.13), we first introduce a crucial decomposition (1.14) and (1.15), which separates the confluence caused by the roughness $F$ from the mathematical difficulties arising from nonlinear terms.
For the homogeneous problem (1.14), the key ingredient is to solve the following linear problem
(1.17) $\left\{\begin{aligned} &y\partial_{x}f_{e}-\partial_{y}^{2}f_{e}=h,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}f_{e}\big|_{y=0}=0,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Taking the Fourier transformation in $x$, the linear operator of the first equation in (1.17) is expressed as $i\xi y-\partial_{y}^{2}$. It shares a structural similarity with the classical Airy operator $\frac{d^{2}}{dz^{2}}-z$. Indeed, there exists a solution $\hat{f}_{e}$ to the following linear problem
$\left\{\begin{aligned} &(i\xi y-\partial_{y}^{2})\hat{f}=\hat{h},\quad(\xi,y)\in\Omega_{M},\\ &\partial_{y}\hat{f}\big|_{y=0}=0,\quad\xi\in{\mathbb{R}}.\end{aligned}\right$.
Similar to the proof in [9], the solution satisfies
${\big\|f_{e}\big\|}_{Y_{0}}\leq C{\big\|h\big\|}_{L^{2}}.$
However, for the existence of the solution, we need to establish the following $L^{2}$ boundedness
${\big\|f_{e}\big\|}_{L^{2}}\lesssim{\big\|h\big\|}_{L^{2}},$
which is the key difference and difficulty compared with the work [9]. We observe that the homogeneous solution to the linear operator involves a factor $(i\xi)^{-\frac{1}{3}}$, which exhibits a singularity near $\xi=0$. For any inhomogeneous term $h$, the general solution includes a term of the form
$(i\xi)^{-\frac{1}{3}}\int G(y,z;\xi)\hat{h}(\xi,z)\,dz,$
where the quantity $(i\xi)^{-\frac{1}{3}}G(y,z;\xi)$ represents a specific Green’s function. Note that the nonlinear term $u\partial_{x}\omega-v\partial_{y}\omega$ cannot be entirely recast solely as $x$-directional derivatives (i.e., $\partial_{x}(\text{Terms})$). This structural constraint hinders low-order estimates derived from the nonlinear terms. To address the singularity near $\xi=0$, a key observation is that the function
$t(z)=\frac{Ai(z)}{Bi(z)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}$
has no zeros in the sector $\{z\in{\mathbb{C}}|\ \text{Arg}(z)\in(-\frac{\pi}{3},\frac{\pi}{3})\}$, where $Ai(z),Bi(z)$ denote two independent Airy functions. This property is non-trivial (see Remark 3.1), which involves the distribution of zeros of the Bessel function. Based on this property, we find a subtle function $G$ to create $(i\xi)^{\frac{1}{3}}$, which takes the following form
$G(y,z;\xi)=\frac{\pi}{t((i\xi)^{\frac{1}{3}}M)}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M,\end{aligned}\right$.
$\mu_{M}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y),$
$\mu_{0}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}Bi((i\xi)^{\frac{1}{3}}y).$
It is easy to verify that $\mu_{M}$ eliminates the singularity near $\xi=0$ and $G$ is bounded on $\big\{z\in{\mathbb{C}}|\ \text{Arg}(z)\in[-\frac{\pi}{6},\frac{\pi}{6}]\big\}$.
The last key point is to show that the displacement $A$ is bounded for all $M\geq M_{0}$. We observe that $A=\int_{0}^{M}\omega(x,y)dy$ and then
${\big\|A\big\|}_{L^{\infty}({\mathbb{R}})}\leq{\big\|\omega\big\|}_{L_{x}^{\infty}L_{y}^{1}}.$
Indeed, we can bound the $L_{x}^{\infty}L_{y}^{1}$ norm of $\omega$ by means of the norm ${\big\|\omega\big\|}_{L^{2}\cap Y_{\bar{\alpha}}}$ for $\bar{\alpha}\geq 0$ via interpolation: there exists some $\theta_{1}=\theta_{1}(\bar{\alpha})\in(0,1)$,
${\big\|\omega\big\|}_{L_{x}^{\infty}L_{y}^{1}}\leq C(\bar{\alpha},M){\big\|\omega\big\|}_{L^{2}}^{\theta_{1}}{\big\|\omega\big\|}_{Y_{0}}^{1-\theta_{1}}+{\big\|\omega\big\|}_{L_{x}^{\infty}({\mathbb{R}};L_{y}^{1}([0,1])}\leq C_{0}(\bar{\alpha},M){\big\|F\big\|}_{H^{2}({\mathbb{R}})}.$
Here $\bar{\alpha}\in[0,\frac{7}{3}]$ and $M\geq M_{0}$. We observe that
(1.18) $\inf_{\bar{\alpha}\in[0,\frac{7}{3}]}C_{0}(\bar{\alpha},M)=C_{0}(0,M)=O([\log(M)]^{\frac{1}{2}}),\quad\forall M\geq M_{0}.$
This means that the bound of $A$ cannot be controlled uniformly in $M$ by the norm ${\big\|F\big\|}_{H^{2}({\mathbb{R}})}$—a contradiction to the physical intuition that that the linear component $(\omega_{0},A_{0})$ of the solution $(\omega,A)$ dictates the behavior of $(\omega,A)$. To eliminate the dependence of the constant $C_{0}$ on $M$, we introduce a new norm ${\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big\|}_{L^{2}}$, which allows us to derive a refined estimate. Specifically, we have
${\big\|\omega\big\|}_{L_{x}^{\infty}L_{y}^{1}}\leq C\big({\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big\|}_{L^{2}}+{\big\|\omega\big\|}_{Y_{0}}\big)\leq C{\big\|F\big\|}_{H^{2}({\mathbb{R}})}.$
Here $C>0$ is independent of $M$. The estimate for this new norm is established by leveraging additional properties of Airy functions (see Proposition 2.1).
1.4. Notations and Organization of the paper
We adopt standard mixed-norms on $\Omega_{M}$ as follows
$\langle f,g\rangle_{L^{2}(\Omega_{M})}=\int_{\Omega_{M}}f\bar{g}\ d{\mathbf{x}},\quad{\big\|f\big\|}_{L^{2}}^{2}=\int_{\Omega_{M}}|f|^{2}d{\mathbf{x}},$
${\big\|f\big\|}_{L_{x}^{p}L_{y}^{q}}={\big\|f\big\|}_{L_{x}^{p}({\mathbb{R}};L_{y}^{q}([0,M]))},\quad\mbox{for }\ p,q\geq 1,$
${\big\|f\big\|}_{H_{x}^{s}L_{y}^{2}}={\big\|f\big\|}_{H_{x}^{s}({\mathbb{R}};L_{y}^{2}([0,M]))},\quad{\big\|f\big\|}_{L_{x}^{2}H_{y}^{s}}={\big\|f\big\|}_{L_{x}^{2}({\mathbb{R}};H_{y}^{s}([0,M]))},\quad\mbox{for }\ s\in{\mathbb{R}}.$
For some inequalities, the notation $B_{0}\lesssim B_{1}$ denotes that there exists some $C>0$ independent of $B_{0}$ and $B_{1}$ such that $B_{0}\leq CB_{1}$.
This paper is organized as follows. First, in Section 2, we characterize the functional spaces used in the analysis of the triple-deck equations and derive key nonlinear estimates, including weighted Sobolev norms for vorticity and displacement. Subsequently, Section 3 focuses on the solvability of the linear system of the triple-deck system. Thereafter, in Section 4, we prove the main result (Theorem 1.2) via a Picard iteration scheme. Finally, we present an appendix, which provides a formal derivation of the complete triple-deck system, including asymptotic matching conditions between the viscous sublayer, main layer, and upper layer.
2. Functional spaces and nonlinear estimates
To study the quantity $\omega$, we first consider the space $X_{0}\cap X_{\alpha}$. Note that the space $X_{\bar{\alpha}}$ consists of a low-order part and a high-order part. The low-order part is characterized by the norms ${\big\|f\big\|}_{L^{2}(\Omega_{M})}$ and ${\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big\|}_{L^{2}(\Omega_{M})}$, while the high-order part is characterized by ${\big\|f\big\|}_{Y_{\bar{\alpha}}}$. We now present some properties between these two parts for $f\in X_{0}$.
Lemma 2.1.
It holds that for any $0\leq\varepsilon\leq\frac{1}{6}$,
${\big\|f\big\|}_{L_{x}^{2}L_{y}^{\infty}}+{\big\||\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}+{\big\|\partial_{y}f\big\|}_{L_{x}^{6}L_{y}^{2}}+{\big\|f\big\|}_{L_{y}^{1}L_{x}^{\infty}}\leq C{\big\|f\big\|}_{X_{0}}.$
Here $C$ is independent of $\varepsilon$.
Proof.
By Sobelev embedding theorem and the interpolation, we obtain
${\big\|f\big\|}_{L_{x}^{2}L_{y}^{\infty}}\leq C{\big\|f\big\|}_{L_{x}^{2}H_{y}^{2}}\leq C{\big\|f\big\|}_{X_{0}},$
${\big\||\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}\leq{\big\||\partial_{x}|^{\frac{1}{3}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}^{6\varepsilon}{\big\||\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}^{1-6\varepsilon}\leq{\big\|f\big\|}_{Y_{0}}.$
Using the inequality ${\big\||\partial_{x}|^{-\frac{1}{3}}f\big\|}_{L_{x}^{6}}\leq C{\big\|f\big\|}_{L_{x}^{2}}$, we derive
${\big\|\partial_{y}f\big\|}_{L_{x}^{6}L_{y}^{2}}\leq C{\big\||\partial_{x}|^{\frac{1}{3}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}.$
Having established the estimates above, we now turn to the $L_{y}^{1}L_{x}^{\infty}$-estimate of $f$. Through direct calculations, we obtain
${\big\|f\big\|}_{L_{x}^{\infty}({\mathbb{R}})}\leq C\min\bigg\{{\big\|y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big\|}_{L_{x}^{2}({\mathbb{R}})}^{\frac{9}{17}}{\big\|y\partial_{x}f\big\|}_{L_{x}^{2}}^{\frac{8}{17}}\cdot y^{-\frac{19}{34}},$
$\qquad\qquad\qquad\qquad\quad\ {\big\|y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big\|}_{L_{x}^{2}({\mathbb{R}})}^{\frac{3}{11}}{\big\||\partial_{x}|^{\frac{2}{3}}f\big\|}_{L_{x}^{2}}^{\frac{8}{11}}\cdot y^{-\frac{1}{22}}\bigg\},$
${\big\|f\big\|}_{L_{y}^{1}L_{x}^{\infty}}\leq C{\big\|y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{3}{11}}{\big\||\partial_{x}|^{\frac{2}{3}}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{8}{11}}{\big\|y^{-\frac{1}{22}}\big\|}_{L_{y}^{2}([0,1])}$
$\qquad\qquad\qquad+C{\big\|y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{9}{17}}{\big\|y\partial_{x}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{8}{17}}{\big\|y^{-\frac{19}{34}}\big\|}_{L_{y}^{2}([1,\infty])}$
$\qquad\qquad\leq C{\big\|y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{3}{11}}{\big\||\partial_{x}|^{\frac{2}{3}}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{8}{11}}$
$\qquad\qquad\qquad+C{\big\|y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{9}{17}}{\big\|y\partial_{x}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{8}{17}}$
$\qquad\qquad\leq C{\big\|f\big\|}_{X_{0}}.$
Lemma 2.2.
Assume that $f\in X_{0}\cap X_{\alpha}$ for $\alpha>0$. Then for any $\varepsilon\in(0,\min\{\frac{1}{6},\frac{\alpha}{2}\}]$, we have
${\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}f\big\|}_{L_{y}^{2}L_{x}^{\infty}}+{\big\|f\big\|}_{L^{\infty}(\Omega_{M})}\leq C\Big(\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}+\frac{1}{\sqrt{\varepsilon}}\Big)\big({\big\|f\big\|}_{X_{0}}+{\big\|f\big\|}_{X_{\alpha}}\big).$
Here $C>0$ is independent of $\varepsilon,\alpha$.
Proof.
By the interpolation and Lemma 2.1, we have
${\big\|\partial_{y}f\big\|}_{L_{x}^{\infty}({\mathbb{R}})}\leq\frac{C}{\sqrt{\varepsilon}}{\big\||\partial_{x}|^{\frac{1}{2}-\varepsilon}\partial_{y}f\big\|}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}}{\big\||\partial_{x}|^{\frac{1}{2}+\varepsilon}\partial_{y}f\big\|}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}},$
${\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}f\big\|}_{L_{y}^{2}L_{x}^{\infty}}\leq\frac{C}{\sqrt{\varepsilon}}{\big\||\partial_{x}|^{\frac{1}{2}+\varepsilon}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L_{y}^{2}L_{x}^{2}}^{\frac{1}{2}}{\big\||\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big\|}_{L_{y}^{2}L_{x}^{2}}^{\frac{1}{2}}$
$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}\Big({\big\||\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}+{\big\||\partial_{x}|^{\frac{1+\alpha}{2}}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}\Big){\big\|f\big\|}_{Y_{0}}^{\frac{1}{2}}$
$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}\big({\big\|f\big\|}_{Y_{0}}+{\big\|f\big\|}_{Y_{\alpha}}\big).$
We denote by $\tilde{f}$ the extension of $f$ to ${\mathbb{R}}^{2}$ such that $\tilde{f}\in H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))\cap H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))$ for $\alpha>0$ and
${\big\|\tilde{f}\big\|}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}\leq C{\big\|f\big\|}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))},$
${\big\|\tilde{f}\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\leq C{\big\|f\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}(0,M))}.$
Direct calculations show that
${\big\|\tilde{f}\big\|}_{L_{x,y}^{\infty}}\leq$ $\ {\big\|\hat{\tilde{f}}\big\|}_{L_{\xi,\eta}^{1}}$
$=$ $\ \int_{{\mathbb{R}}^{2}}\bigg\{\big|\hat{\tilde{f}}(\xi,\eta)(1+|\eta|)^{2}\big|^{\theta}\big|\hat{\tilde{f}}(\xi,\eta)(1+|\xi|)^{\frac{2}{3}+\frac{\alpha}{2}}\big|^{1-\theta}$
$\qquad\quad\cdot(1+|\eta|)^{-2\theta}(1+|\xi|)^{-(\frac{2}{3}+\frac{\alpha}{2})(1-\theta)}\bigg\}d\xi d\eta$
$\leq$ $\ C(\alpha,\theta){\big\|\big|\hat{\tilde{f}}(\xi,\eta)(1+|\eta|)^{2}\big|^{\theta}\big\|}_{L_{\xi,\eta}^{\frac{2}{\theta}}}{\big\|\big|\hat{\tilde{f}}(\xi,\eta)(1+|\xi|)^{\frac{2}{3}+\frac{\alpha}{2}}\big|^{1-\theta}\big\|}_{L_{\xi,\eta}^{\frac{2}{1-\theta}}}$
$\leq$ $\ C(\alpha,\theta){\big\|\tilde{f}\big\|}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}^{\theta}{\big\|\tilde{f}\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}^{1-\theta}$
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|\tilde{f}\big\|}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big\|\tilde{f}\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\big).$
Here $C(\alpha,\theta)={\big\|(1+|\eta|)^{-2\theta}(1+|\xi|)^{-(\frac{2}{3}+\frac{\alpha}{2})(1-\theta)}\big\|}_{L_{\xi,\eta}^{2}}$ and $\theta=\frac{1}{2}(\frac{1}{4}+\frac{1+3\alpha}{4+3\alpha})$. This shows that
${\big\|f\big\|}_{L_{x,y}^{\infty}}\leq$ $\ {\big\|\tilde{f}\big\|}_{L_{x,y}^{\infty}}\leq C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|\tilde{f}\big\|}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big\|\tilde{f}\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\big)$
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|f\big\|}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))}+{\big\|f\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}\big({\mathbb{R}};L_{y}^{2}((0,M))\big)}\big)$
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|f\big\|}_{X_{0}}+{\big\|f\big\|}_{X_{\alpha}}\big).$
Next we are devoted to the estimates of nonlinear terms. Let the velocity $(u,v)$ satisfy the divergence-free condition
$\partial_{x}u+\partial_{y}v=0$
subject to the boundary conditions
$(u,v)\big|_{y=0}=0,\quad u\big|_{y=M}=A_{M}.$
where $A_{M}$ denotes a fixed velocity at the upper boundary $y=M$. Define the function
$h=-(u\partial_{x}\omega+v\partial_{y}\omega),$
for $\omega=\partial_{y}u$. Using the divergence-free condition and boundary conditions, we can rewrite $h$ as
$h=-u\partial_{x}\omega-(V^{0}-y\partial_{x}A_{M})\partial_{y}\omega.$
Here
(2.1) $V^{0}=\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz.$
Proposition 2.1.
Let $\alpha>\frac{2}{3},\ \varepsilon\in(0,\frac{1}{6}]$. Assume that $\omega\in X_{0}\cap X_{\alpha}$ and $\partial_{x}A_{M}\in L^{\infty}\cap H^{\frac{\alpha}{2}}$. Then $h\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$ with
(2.2) ${\big\|h\big\|}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}\big){\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}},$
${\big\||\partial_{x}|^{\frac{\alpha}{2}}h\big\|}_{L^{2}(\Omega_{M})}\leq CM\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\big){\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}.$
Here $C=C(\alpha,\varepsilon)>0$ is independent of $M$ and $A_{M}$.
Proof.
By virtue of Lemma 2.1 and Lemma 2.2, we obtain
${\big\|h\big\|}_{L^{2}(\Omega_{M})}\leq$ $\ C{\big\|\frac{u}{y}\big\|}_{L^{\infty}(\Omega_{M})}{\big\|y\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}+CM^{\frac{3\varepsilon}{2}}{\big\|\tilde{z}\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}$
$\ \ +CM^{\frac{1+3\varepsilon}{2}}{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}$
$\leq$ $\ C\big({\big\|\omega\big\|}_{L^{\infty}(\Omega_{M})}+M^{\frac{3\varepsilon}{2}}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}\big){\big\|y\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}$
$\ \ +CM^{\frac{1+3\varepsilon}{2}}{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}$
$\leq$ $\ C(1+M^{\frac{1+3\varepsilon}{2}}){\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2}+CM^{\frac{1+3\varepsilon}{2}}{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}},$
which gives the first estimate of (2.2).
For the higher-order estimates, by product estimates, we have
(2.3) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big\|}_{L^{2}(\Omega_{M})}$
$\leq C{\big\|\omega\big\|}_{L^{\infty}(\Omega_{M})}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}+C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big\|y|\partial_{x}|^{\frac{1}{3}}\omega\big\|}_{L^{\infty}(\Omega_{M})}$
$\leq C{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2}+C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big\|y|\partial_{x}|^{\frac{1}{3}}\omega\big\|}_{L^{\infty}(\Omega_{M})},$
and
(2.4) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(V^{0}\partial_{y}\omega)\big\|}_{L^{2}(\Omega_{M})}$
$\leq C{\big\|y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}|\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big\|}_{L_{x}^{2}L_{y}^{\infty}}{\big\|y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{\infty}L_{y}^{2}}$
$\qquad+C{\big\|y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big\|}_{L_{x}^{4}L_{y}^{\infty}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}$
$\leq CM^{\frac{3\varepsilon}{2}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}\tilde{z}\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}$
$\qquad+C{\big\|\tilde{z}\partial_{x}\omega\big\|}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}$
$\leq CM^{\frac{3\varepsilon}{2}}{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2}+C{\big\|\tilde{z}\partial_{x}\omega\big\|}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}.$
It remains to estimate the right-hand side of the above inequality. By the Sobolev embedding, we have
(2.5) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}\leq C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+C{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{2}}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}$
$\qquad\leq C{\big\|\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{3}}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}\leq C{\big\|\omega\big\|}_{X_{\alpha}},$
${\big\|y|\partial_{x}|^{\frac{1}{3}}\omega\big\|}_{L^{\infty}(\Omega_{M})}\leq\ M{\big\||\partial_{x}|^{\frac{1}{3}}\omega\big\|}_{L^{\infty}(\Omega)}$
$\qquad\leq C\frac{[1+(\alpha-\frac{2}{3})]^{\frac{1}{2}}}{\alpha-\frac{2}{3}}M{\big\||\partial_{x}|^{\frac{1}{3}}\omega\big\|}_{H_{x}^{\frac{2}{3}+(\frac{\alpha}{2}-\frac{1}{3})}L_{y}^{2}\cap L_{x}^{2}H_{y}^{2}}\quad\mbox{for}\ \ \alpha>\frac{2}{3}$
$\qquad\leq CM\big({\big\||\partial_{x}|^{\frac{1}{3}}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{2}}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{1}{3}}\partial_{y}^{2}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}\big)$
$\qquad\leq CM\big({\big\|\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{2}}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\|\partial_{y}^{2}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{\alpha}{2}}\partial_{y}^{2}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}\big)$
$\qquad\leq CM{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}},$
and
(2.6) ${\big\|y\partial_{x}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}$ $\leq C{\big\||\partial_{x}|^{\frac{1}{4}}y\partial_{x}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}$
$\leq C\big({\big\|y\partial_{x}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}\big)\quad\mbox{for}\ \ \alpha\geq\frac{1}{2}$
$\leq C{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}},$
${\big\||\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}$ $\leq{\big\||\partial_{x}|^{\frac{\alpha}{2}+\frac{1}{4}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}$
$\leq C\big({\big\||\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{\alpha}{2}+\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}\big)$
$\leq C{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}.$
Combining all estimates above, we deduce that for $\alpha>\frac{2}{3}$,
(2.7) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big\|}_{L^{2}(\Omega_{M})}\leq\ CM{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2},$
${\big\||\partial_{x}|^{\frac{\alpha}{2}}(V^{0}\partial_{y}\omega)\big\|}_{L^{2}(\Omega_{M})}\leq\ CM^{\frac{3\varepsilon}{2}}{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2}.$
Moreover, we have
(2.8) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(y\partial_{x}A_{M}\partial_{y}\omega)\big\|}_{L^{2}(\Omega_{M})}$
$\leq C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}{\big\|y\partial_{y}\omega\big\|}_{L_{x}^{\infty}L_{y}^{2}}+C{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}({\mathbb{R}})}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y\partial_{y}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}$
$\leq CM^{\frac{1+3\varepsilon}{2}}{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}{\big\|\partial_{x}A_{M}\big\|}_{\dot{H}^{\frac{\alpha}{2}}}+CM{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}}$
$\leq CM{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}.$
Substituting (2.5)-(2.8) into (2.3)-(2.4), we conclude the last estimate of (2.2). ∎
Remark 2.1.
Indeed, if $h$ is replaced by $h_{1}=-(u\partial_{x}\omega_{1}+v\partial_{y}\omega_{1})$ for $\omega_{1}\in X_{0}\cap X_{\alpha}$, then by repeating the proof of Proposition 2.1, we have
${\big\|h_{1}\big\|}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}\big){\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}},\ \ \varepsilon\in(0,\frac{1}{6}],$
${\big\||\partial_{x}|^{\frac{\alpha}{2}}h_{1}\big\|}_{L^{2}(\Omega_{M})}\leq CM\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\big){\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}.$
The following lemma is devoted to the integral boundary condition $I_{M}[\omega]=A(x)+F(x)$.
Lemma 2.3.
It holds that for any $\bar{\alpha}\geq 0$,
(2.9) ${\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big\|}_{L^{2}({\mathbb{R}})}\leq$ $\ C{\big\|f\big\|}_{Y_{\bar{\alpha}}}.$
Here $C$ is independent of $\bar{\alpha}$.
Proof.
Observe that
$\left|\int_{0}^{M}\hat{f}(\xi,y)dy\right|=$ $\left|\int_{0}^{|\xi|^{-\frac{1}{3}}}\hat{f}(\xi,y)dy+\int_{|\xi|^{-\frac{1}{3}}}^{M}y^{-1}y\hat{f}(\xi,y)dy\right|$
$\leq$ $\ C|\xi|^{-\frac{1}{6}}{\big\|\hat{f}(\xi,\cdot)\big\|}_{L_{y}^{2}}+C|\xi|^{\frac{1}{6}}{\big\|y\hat{f}(\xi,\cdot)\big\|}_{L_{y}^{2}},$
which implies
$\big|\xi^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big|\leq$ $\ C|\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\cdot{\big\|\hat{f}(\xi,\cdot)\big\|}_{L_{y}^{2}({\mathbb{R}})}+C|\xi|^{1+\frac{\bar{\alpha}}{2}}{\big\|y\hat{f}(\xi,\cdot)\big\|}_{L_{y}^{2}}).$
Thus, we obtain
${\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big\|}_{L^{2}({\mathbb{R}})}\leq C\big({\big\||\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big\|}_{L^{2}}+{\big\|y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}f\big\|}_{L^{2}}\big)\leq C{\big\|f\big\|}_{Y_{\bar{\alpha}}}.$
It completes the proof. ∎
3. The solvability of the linear system
Recall that, to solve the nonlinear system (1.13), we decompose $\omega=\omega_{0}+\bar{\omega}$ and $A=A_{0}+\bar{A}$ with
(3.1) $\left\{\begin{aligned} &y\partial_{x}\omega_{0}-\partial_{y}^{2}\omega_{0}=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\omega_{0}\big|_{y=0}=\partial_{x}|\partial_{x}|A_{0},\quad I_{M}[\omega_{0}]=A_{0}+F,\quad x\in{\mathbb{R}},\end{aligned}\right$.
and
(3.2) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}-\partial_{y}^{2}\bar{\omega}=-(u\partial_{x}\omega+v\partial_{y}\omega)\triangleq{\bar{h}},\quad(x,y)\in\Omega_{M},\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\bar{\omega}\big|_{y=0}=\partial_{x}|\partial_{x}|\bar{A},\quad I_{M}[\bar{\omega}]=\bar{A},\quad x\in{\mathbb{R}}.\end{aligned}\right$.
To address the system (3.2), we first solve an inhomogeneous linear problem without the integral condition
(3.3) $\left\{\begin{aligned} &y\partial_{x}f_{e}-\partial_{y}^{2}f_{e}=h_{e},\quad(x,y)\in\Omega_{M},\\ &\partial_{y}f_{e}\big|_{y=0}=0,\quad f_{e}\big|_{y=M}=0,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Proposition 3.1.
Assume that $h_{e}\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$ for some $\alpha>0$. Then there exists a unique solution $f_{e}\in X_{0}\cap X_{\alpha}$ to the problem (3.3). Moreover, there holds that
${\big\|f_{e}\big\|}_{X_{0}\cap X_{\alpha}}\leq C\big[(1+M^{2}){\big\|h_{e}\big\|}_{L^{2}(\Omega_{M})}+{\big\||\partial_{x}|^{\frac{\alpha}{2}}h_{e}\big\|}_{L^{2}(\Omega_{M})}\big],$
where $C$ is a constant independent of $\alpha,M$.
Proof.
The proof proceeds in three steps.
Step 1. The explicit formula of the solution.
Taking the Fourier transform in variable $x$ into (3.3), we obtain
(3.4) $\left\{\begin{aligned} &i\xi y\hat{f}_{e}-\partial_{y}^{2}\hat{f}_{e}=\hat{h}_{e},\quad(\xi,y)\in{\mathbb{R}}\times(0,M),\\ &\partial_{y}\hat{f}_{e}\big|_{y=0}=0,\ \hat{f}_{e}\big|_{y=M}=0,\quad\xi\in{\mathbb{R}}.\end{aligned}\right$.
Using the Green-function method, we find the unique solution $\hat{f}_{e}$ of (3.4), which has an explicit formula
(3.5) $\hat{f}_{e}(\xi,y)=$ $\ (i\xi)^{-\frac{1}{3}}\int_{0}^{M}G(y,z;\xi)\hat{h}_{e}(\xi,z)dz,$
with the definition of $G$ given by
(3.6) $G(y,z;\xi)=\pi s_{M}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M,\end{aligned}\right$.
$s_{M}=\bigg(\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}\bigg)^{-1},$
$\mu_{M}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y),$
$\mu_{0}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}Bi((i\xi)^{\frac{1}{3}}y).$
Step 2. The low-order estimates of $f_{e}$.
To this end, we first claim that $s_{M}$ is bounded. Since the ratio $\frac{Ai^{\prime}(z)}{Bi^{\prime}(z)}$ of the Airy functions at $z=0$ is $\frac{1}{\sqrt{3}}$, we then define
$t(z):=\frac{Ai(z)}{Bi(z)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}=\frac{Ai(z)}{Bi(z)}+\frac{1}{\sqrt{3}}=\frac{\sqrt{3}Ai(z)+Bi(z)}{\sqrt{3}Bi(z)}.$
It is well-known that $Ai(z)$ and $Bi(z)$ are holomorphic on $\mathbb{C}$} and do not vanish for $\text{Arg}(z)\in(-\frac{\pi}{3},\frac{\pi}{3})$ [28, Page 15]. For clarity of the proof, we claim that $\sqrt{3}Ai(z)+Bi(z)$ on $\mathbb{C}$} has all its zeros in the rays $\big\{z\in{\mathbb{C}}|\ \text{Arg}(z)=\frac{\pi}{3}+\frac{2k\pi}{3},\ \ k\in{\mathbb{Z}}\big\}$ (see details in Remark 3.1). Define a sector with the angle $2\theta$ by
$C_{\theta}=\big\{z\in{\mathbb{C}}\big|\ \text{Arg}(z)\in(-\theta,\theta)\big\}.$
Then, $t(z)$ is a holomorphic and non-vanishing function on $C_{\frac{\pi}{3}}$. Moreover, for any $\delta\in(0,\frac{\pi}{3})$, the Airy functions $Ai,Bi$ satisfy the following asymptotic expansions(see [20, Chapter 9]):
(3.7) $Ai(z)\sim\frac{1}{3^{\frac{2}{3}}\Gamma(\frac{1}{3})}-\frac{1}{3^{\frac{1}{3}}\Gamma(\frac{2}{3})}z,\quad Bi(z)\sim\frac{1}{3^{\frac{1}{6}}\Gamma(\frac{1}{3})}+\frac{3^{\frac{1}{6}}}{\Gamma(\frac{2}{3})}z,\qquad\mbox{as}\ \ z\to 0,$
$Ai(z)\sim\frac{e^{-\frac{2}{3}z^{\frac{3}{2}}}}{2\sqrt{\pi}z^{\frac{1}{4}}},\quad Bi(z)\sim\frac{e^{\frac{2}{3}z^{\frac{3}{2}}}}{\sqrt{\pi}z^{\frac{1}{4}}},\qquad\mbox{as}\ \ |z|\to\infty\ \ \big(z\in C_{\frac{\pi}{3}-\delta}\big).$
By the asymptotic behavior (3.7), we have
(3.8) $\lim_{|z|\to\infty,\atop|\text{Arg}(z)|\in[0,\frac{\pi}{3}-\delta)}t(z)=\frac{1}{\sqrt{3}}.$
Then, combining the holomorphicity and non-vanishing property of $t(z)$ on $C_{\frac{\pi}{3}}$ with (3.8), there exists some $c_{0}>0$ such that
$|t(z)|\geq c_{0}^{-1},\quad\text{for }z\in\overline{C_{\frac{\pi}{6}}}.$
Thanks to $s_{M}=\bigg(\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}\bigg)^{-1}=\frac{1}{t((i\xi)^{\frac{1}{3}}M)}$, we then obtain
(3.9) $|s_{M}|\leq c_{0},\qquad\forall\ \xi\in{\mathbb{R}},\ M>0.$
To derive a low-order estimate of $f_{e}$, we decompose the frequency space into two regimes. For the low-frequency regime where $|\xi|\leq M^{-3}$, we observe that $|(i\xi)^{\frac{1}{3}}y|\leq 1$ for all $y\in[0,M]$. By the continuity of $Ai$ and $Bi$ on $\mathbb{C}$}, we have
(3.10) $|\mu_{0}(\xi,y)|=\big|Ai((i\xi)^{\frac{1}{3}}y)+\frac{1}{\sqrt{3}}Bi((i\xi)^{\frac{1}{3}}y)\big|\leq C,\quad\forall\ y\in[0,M].$
Next, we compute
(3.11) $\mu_{M}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y)$
$\qquad\quad\ \ =Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai\big((i\xi)^{\frac{1}{3}}(M-y)+(i\xi)^{\frac{1}{3}}y\big)}{Bi\big((i\xi)^{\frac{1}{3}}(M-y)+(i\xi)^{\frac{1}{3}}y\big)}Bi((i\xi)^{\frac{1}{3}}y).$
Since $Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Bi((i\xi)^{\frac{1}{3}}y)}Bi((i\xi)^{\frac{1}{3}}y)=0$ and $Bi(z)\not=0$ for ${\rm Arg}(z)\in(-\frac{\pi}{3},\frac{\pi}{3})$, an application of the Mean Value Theorem to (3.11) yields
(3.12) $|\mu_{M}(\xi,y)|\leq C|\xi|^{\frac{1}{3}}(M-y),\quad\forall\ y\in[0,M].$
It follows that
(3.13) $|\hat{f}_{e}(\xi,y)|=$ $\ \bigg|\pi s_{M}(i\xi)^{-\frac{1}{3}}\bigg\{\mu_{M}(\xi,y)\cdot\int_{0}^{y}\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)dz$
$\qquad\qquad\qquad\quad+\mu_{0}(\xi,y)\cdot\int_{y}^{M}\mu_{M}(\xi,z)\hat{h}_{e}(\xi,z)dz\bigg\}\bigg|$
$\leq$ $\ C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg\{\big|\mu_{M}(\xi,y)\big|\cdot\int_{0}^{y}\big|\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big|dz$
$\qquad\qquad\qquad\quad+\big|\mu_{0}(\xi,y)\big|\cdot\int_{y}^{M}\big|\mu_{M}(\xi,z)\hat{h}_{e}(\xi,z)\big|dz\bigg\}$
[…]
$\quad+C|\xi|^{-\frac{1}{3}}{\big\||\xi|^{\frac{1}{3}}(M-z)\big\|}_{L_{z}^{2}([y,M])}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}([y,M])}$
$\leq$ $\ C(M-y)M^{\frac{1}{2}}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}([0,M])},\quad\mbox{for}\ \ y\in[0,M].$
Hence, we obtain a point-wise estimate of $\hat{f}_{e}$ for the low frequency as follows
(3.14) ${\big\|(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\hat{f}_{e}(\xi,\cdot)\big\|}_{L_{y}^{2}(0,M)}\leq$ $\ CM^{2}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}(0,M)},$
for $\alpha^{\prime}\geq 0$.
For the high frequency $|\xi|\geq M^{-3}$, by the asymptotic behavior (3.7) of Airy functions, we have
(3.15) $|\mu_{M}(\xi,y)|=\big|Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y)\big|$
$\qquad\qquad\ \leq C[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{4}}e^{-\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}},$
$|\mu_{0}(\xi,y)|=\big|Ai((i\xi)^{\frac{1}{3}}y)+\frac{1}{\sqrt{3}}Bi((i\xi)^{\frac{1}{3}}y)\big|$
$\qquad\qquad\leq C[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{4}}\bigg[e^{-\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}}+e^{\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}}\bigg]$
$\qquad\qquad\leq C[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{4}}e^{\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}},$
and then,
(3.16) ${\big\|\mu_{0}(\xi,z)\big\|}_{L_{z}^{2}(0,y)}\leq C|\xi|^{-\frac{1}{6}}\bigg({\big\|Ai(z)\big\|}_{L_{z}^{2}(0,(i\xi)^{\frac{1}{3}}y)}+{\big\|Bi(z)\big\|}_{L_{z}^{2}(0,(i\xi)^{\frac{1}{3}}y)}\bigg)$
$\qquad\qquad\qquad\quad\leq C|\xi|^{-\frac{1}{6}}[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{2}}e^{\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}},$
${\big\|\mu_{M}(\xi,z)\big\|}_{L_{z}^{2}(y,M)}\leq C{\big\|[1+|\xi|^{\frac{1}{3}}z]^{-\frac{1}{4}}e^{-\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}z]^{\frac{3}{2}}}\big\|}_{L_{z}^{2}(y,M)}$
$\qquad\qquad\qquad\qquad\leq C|\xi|^{-\frac{1}{6}}{\big\|[1+\tilde{z}]^{-\frac{1}{4}}e^{-\frac{2}{3}\tilde{z}^{\frac{3}{2}}}\big\|}_{L_{\tilde{z}}^{2}(({\rm Re\ }[(i\xi)^{\frac{1}{3}}y],M))}$
$\qquad\qquad\qquad\qquad\leq C|\xi|^{-\frac{1}{6}}[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{2}}e^{-\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}}.$
Combining (3.5), (3.6), (3.9), (3.15) with (3.16), we obtain
$|\hat{f}_{e}(\xi,y)|\leq C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg\{\big|\mu_{M}(\xi,y)\big|\cdot\int_{0}^{y}\big|\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big|dz$
$\qquad\qquad\qquad\qquad\qquad\quad+\big|\mu_{0}(\xi,y)\big|\cdot\int_{y}^{M}\big|\mu_{M}(\xi,z)\hat{h}_{e}(\xi,z)\big|dz\bigg\}$
$\qquad\leq C|\xi|^{-\frac{1}{3}}[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{4}}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}([0,M])}$
$\qquad\qquad\cdot\bigg\{e^{-\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}}\cdot{\big\|\mu_{0}(\xi,z)\big\|}_{L_{z}^{2}([0,y])}$
$\qquad\qquad\quad+e^{\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}}\cdot{\big\|\mu_{M}(\xi,z)\big\|}_{L_{z}^{2}([y,M])}\bigg\}$
$\qquad\leq C|\xi|^{-\frac{1}{2}}[1+|\xi|^{\frac{1}{3}}y]^{-\frac{3}{4}}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}([0,M])},\quad\mbox{for}\ \ y\in[0,M].$
Hence, we obtain a pointwise estimate of $\hat{f}_{e}$ for the high frequency as follows
(3.17) ${\big\|(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\hat{f}_{e}(\xi,\cdot)\big\|}_{L_{y}^{2}(0,M)}\leq C(\alpha^{\prime})|\xi|^{-\frac{1}{2}}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}(0,M)}\cdot|\xi|^{-\frac{1}{6}}$
$\qquad\leq C|\xi|^{-\frac{2}{3}}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}(0,M)}\leq CM^{2}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}(0,M)},$
for $\alpha^{\prime}\in[0,\frac{1}{4})$. By (3.14) and (3.17), we obtain
(3.18) ${\big\|f_{e}\big\|}_{L^{2}(\Omega_{M})}+{\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f_{e}\big\|}_{L^{2}(\Omega_{M})}\leq CM^{2}{\big\|h_{e}\big\|}_{L^{2}(\Omega_{M})}\leq CM^{2}{\big\|h_{e}\big\|}_{L^{2}(\Omega_{M})}.$
Here $C$ is independent of $M$ and $f_{e}$.
Step 3. The high-order estimates of $f_{e}$.
For $\bar{\alpha}\geq 0$, we set
(3.19) $B_{0}^{2}=|\xi|^{2+\bar{\alpha}}{\big\|y\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+|\xi|^{\frac{4}{3}+\bar{\alpha}}{\big\|\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}.$
Multiplying the equation (3.4) by $\xi^{1+\bar{\alpha}}y\hat{f}_{e}$, we obtain
$i|\xi|^{2+\bar{\alpha}}{\big\|y\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}{\big\|y^{\frac{1}{2}}\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}\langle\partial_{y}\hat{f}_{e},\hat{f}_{e}\rangle_{L_{y}^{2}}=\langle\hat{h}_{e},y\xi^{1+\bar{\alpha}}\hat{f}_{e}\rangle_{L_{y}^{2}}.$
Here we used the fact that $\hat{f}_{e}\big|_{y=M}=0$. Set
(3.20) $B_{1}^{2}=$ $\ |\xi|^{1+\bar{\alpha}}{\big\|y^{\frac{1}{2}}\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2},$
$B_{2}^{2}=$ $\ |\xi|^{\frac{5}{3}+\bar{\alpha}}{\big\|y^{\frac{1}{2}}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big\|\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}.$
Then we have
(3.21) $|\xi|^{2+\bar{\alpha}}{\big\|y\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+B_{1}^{2}$ $\leq\big|-\xi^{1+\bar{\alpha}}\langle\partial_{y}\hat{f}_{e},\hat{f}_{e}\rangle_{L_{y}^{2}}+\langle\hat{h}_{e},y\xi^{1+\bar{\alpha}}\hat{f}_{e}\rangle_{L_{y}^{2}}\big|$
$\leq\big|-\langle\xi^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e},\xi^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}+\langle\xi^{\frac{\bar{\alpha}}{2}}\hat{h}_{e},y\xi^{1+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}\big|$
$\leq C\bigg(|\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}{\big\|\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}\bigg)B_{0}$
$\leq CB_{0}(B_{2}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}).$
Here $C$ is independent of $\bar{\alpha}$. Multiplying the equation (3.4) by $|\xi|^{\frac{2}{3}+\bar{\alpha}}\hat{f}_{e}$, we obtain
(3.22) $i\xi|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big\|y^{\frac{1}{2}}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big\|\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}=\langle\hat{h}_{e},|\xi|^{\frac{2}{3}+\bar{\alpha}}\hat{f}_{e}\rangle,\quad\mbox{by}\ \ \hat{f}_{e}(\cdot,M)=0.$
Then we have
$|\xi|^{\frac{5}{3}+\bar{\alpha}}{\big\|y^{\frac{1}{2}}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big\|\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}$ $\leq{\big\||\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\big\|}_{L_{y}^{2}}{\big\||\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big\|}_{L_{y}^{2}}$
$\leq CB_{0}{\big\||\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big\|}_{L_{y}^{2}}.$
Thus, we get by (3.20) that
(3.23) $B_{2}^{2}\leq$ $\ CB_{0}{\big\||\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big\|}_{L_{y}^{2}}.$
On the other hand, we have
(3.24) $|\xi|^{\frac{4}{3}+\bar{\alpha}}{\big\|\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}$ $=|\xi|^{\frac{4}{3}+\bar{\alpha}}\int_{0}^{M}\partial_{y}(y)|\hat{f}_{e}|^{2}dy$
$=|\xi|^{\frac{4}{3}+\bar{\alpha}}M\hat{f}_{e}^{2}\big|_{y=M}-|\xi|^{\frac{4}{3}+\bar{\alpha}}\int_{0}^{M}y(\bar{\hat{f}}_{e}\partial_{y}\hat{f}_{e}+\hat{f}_{e}\partial_{y}\bar{\hat{f}}_{e})dy$
$\leq 2{\big\||\xi|^{1+\frac{\bar{\alpha}}{2}}y\hat{f}_{e}\big\|}_{L_{y}^{2}}{\big\||\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}$
$\leq CB_{0}B_{2}.$
By (3.19), (3.20), (3.21), (3.23) and (3.24), we have
(3.25) $B_{0}^{2}+B_{1}^{2}\leq C|\xi|^{\bar{\alpha}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}^{2}.$
Then by (3.23) and (3.25), we get
(3.26) $B_{2}^{2}\leq$ $\ CB_{0}^{2}+C|\xi|^{\bar{\alpha}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}^{2}\leq C|\xi|^{\bar{\alpha}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}^{2}.$
Applying Plancherel’s theorem, we obtain
(3.27) ${\big\|f_{e}\big\|}_{Y_{\bar{\alpha}}}^{2}\leq$ $\ C\bigg({\big\||\partial_{x}|^{\frac{\bar{\alpha}}{2}}y\partial_{x}f_{e}\big\|}_{L^{2}}^{2}+{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f_{e}\big\|}_{L^{2}}^{2}+{\big\||\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}f_{e}\big\|}_{L^{2}}^{2}$
$\quad+{\big\||\partial_{x}|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}f_{e}\big\|}_{L^{2}}^{2}+{\big\||\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}f_{e}\big\|}_{L^{2}}^{2}\bigg)$
$=$ $\ C\bigg({\big\||\xi|^{1+\frac{\bar{\alpha}}{2}}y\hat{f}_{e}\big\|}_{L^{2}}^{2}+{\big\||\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\big\|}_{L^{2}}^{2}+{\big\||\xi|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}\hat{f}_{e}\big\|}_{L^{2}}^{2}$
$\quad+{\big\||\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e}\big\|}_{L^{2}}^{2}+{\big\||\xi|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}\hat{f}\big\|}_{L^{2}}^{2}\bigg)$
$\leq$ $\ C\bigg(\int_{{\mathbb{R}}}B_{0}^{2}+B_{1}^{2}+B_{2}^{2}+B_{0}^{2}+{\big\||\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big\|}_{L_{y}^{2}}^{2}d\xi\bigg)$
[…]
$=$ $\ C{\big\||\partial_{x}|^{\frac{\bar{\alpha}}{2}}h_{e}\big\|}_{L^{2}}^{2}.$
Finally, it follows from (3.18) and (3.27) that
(3.28) ${\big\|f_{e}\big\|}_{X_{0}\cap X_{\alpha}}\leq C\big[(1+M^{2}){\big\|h_{e}\big\|}_{L^{2}}+{\big\||\partial_{x}|^{\frac{\alpha}{2}}h_{e}\big\|}_{L^{2}}\big].$
Here $C$ is independent of $\alpha,M$, and $f_{e}$. ∎
Remark 3.1.
The Airy functions $Ai(z),Bi(z)$ admit decompositions [20, Chapter 9] as follows
$Ai(z)=\frac{\sqrt{z}}{3}\big(I_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})-I_{\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})\big),\quad z\in{\mathbb{C}},$
$Bi(z)=\sqrt{\frac{z}{3}}\big(I_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})+I_{\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})\big),\quad z\in{\mathbb{C}},$
$I_{\nu}(z)=e^{-\frac{\nu\pi i}{2}}J_{\nu}(ze^{\frac{\pi i}{2}}),\quad-\pi\leq\text{Arg}(z)\leq\frac{\pi}{2},\ \nu\in{\mathbb{R}}.$
Here $J_{\nu}(z)=\sum\limits_{k=0}^{\infty}\frac{(-1)^{k}\big(\frac{z}{2}\big)^{2k+\nu}}{k!\Gamma(k+\nu+1)}$ denotes the Bessel function satisfying the Bessel equation
$\Big(z^{2}\frac{d^{2}}{dz}+z\frac{d}{dz}+(z^{2}-\nu^{2})\Big)J_{\nu}(z)=0.$
We then derive
(3.29) ${\mathcal{C}}(z)$= $\ \sqrt{3}Ai(z)+Bi(z)=\sqrt{\frac{4z}{3}}I_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})$
$=$ $\ \sqrt{\frac{4z}{3}}e^{\frac{\pi i}{6}}J_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}}e^{\frac{\pi i}{2}}),\qquad-\frac{2\pi}{3}\leq\text{Arg}(z)\leq\frac{\pi}{3}.$
For $\nu>-1$, the Bessel function $J_{\nu}(z)$ has only real zeros in $\mathbb{C}$} [31, Page 482]. Then by (3.29), the zeros of ${\mathcal{C}}(z)$ in the sector $\big\{z\in{\mathbb{C}}|\ -\frac{2\pi}{3}\leq\text{Arg}(z)\leq\frac{\pi}{3}\big\}$ lie only on rays $\big\{z\in{\mathbb{C}}|\ \text{Arg}(z)=-\frac{\pi}{3}\ \text{or}\ \frac{\pi}{3}\big\}$. Notice that ${\mathcal{C}}(z)$ is entire and ${\mathcal{C}}(z)={\mathcal{C}}(ze^{\frac{2\pi i}{3}})$ for any $z\in{\mathbb{C}}$, then we have that all zeros of ${\mathcal{C}}(z)$ in $\mathbb{C}$} belong to $\big\{z\in{\mathbb{C}}|\ \text{Arg}(z)=\frac{\pi}{3}+\frac{2k\pi}{3},\ k\in{\mathbb{Z}}\big\}$.
We next solve the following homogeneous problem
(3.30) $\left\{\begin{aligned} &y\partial_{x}\omega_{b}-\partial_{y}^{2}\omega_{b}=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\omega_{b}\big|_{y=0}=\partial_{x}|\partial_{x}|A,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Proposition 3.2.
Let $\bar{\alpha}\geq 0$ and $g=\partial_{x}|\partial_{x}|A\in H^{\frac{1}{6}+\frac{\bar{\alpha}}{2}}\cap\dot{H}^{-1}$. Then there exists a solution $\omega_{b}\in X_{\bar{\alpha}}$ to the boundary-value problem (3.30) with the estimate
(3.31) ${\big\|\omega_{b}\big\|}_{X_{\bar{\alpha}}}\leq C{\big\||\partial_{x}|^{\frac{1}{6}+\frac{\bar{\alpha}}{2}}g\big\|}_{L^{2}({\mathbb{R}})}\leq C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})},$
where $C>0$ is a constant independent of $\bar{\alpha}$ and $A$.
Proof.
To solve the boundary-value problem (3.30), we take the Fourier transform in variable $x$. Then we have
(3.32) $\left\{\begin{aligned} i\xi y\hat{\omega}_{b}-\partial_{y}^{2}\hat{\omega}_{b}=&\ 0,\quad(\xi,y)\in{\mathbb{R}}\times(0,M)\\ \partial_{y}\hat{\omega}_{b}\big|_{y=0}=\hat{g}=&\ i\xi|\xi|\hat{A},\quad\xi\in{\mathbb{R}}.\end{aligned}\right$.
The general solution to the BVP (3.32) is as follows
(3.33) $\hat{\omega}_{b}(\xi,y)=C(\xi)\bigg(\sqrt{3}Ai((i\xi)^{\frac{1}{3}}y)+Bi((i\xi)^{\frac{1}{3}}y)\bigg)+\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}\hat{g}$.
Taking $C(\xi)=0$, we obtain a solution $\omega_{b}$ with the explicit formula
(3.34) $\hat{\omega}_{b}(\xi,y)=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}\hat{g}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}\hat{g}(\xi),\quad\xi<0.\end{aligned}\right$.
Then we have
(3.35) ${\big\|(|\partial_{x}|^{\frac{1}{3}}y)^{\alpha^{\prime}}\omega_{b}\big\|}_{L^{2}(\Omega_{M})}$ $\leq C{\big\|(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\hat{\omega}_{b}\big\|}_{L^{2}(\Omega_{M})}$
$\leq C{\big\|{\big\|Ai((i\xi)^{\frac{1}{3}}y)(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\big\|}_{L_{y}^{2}}\cdot|\xi|^{\frac{5}{3}}|\hat{A}(\xi)|\big\|}_{L_{\xi}^{2}}$
$\leq C{\big\||\xi|^{\frac{3}{2}}\hat{A}(\xi)\big\|}_{L_{\xi}^{2}}=C{\big\||\partial_{x}|^{\frac{3}{2}}A\big\|}_{L^{2}({\mathbb{R}})}$
$\leq C{\big\||\partial_{x}|^{\frac{5}{6}}A\big\|}_{H^{\frac{2}{3}}({\mathbb{R}})}.$
As $g\in H^{\frac{1}{6}+\frac{\alpha}{2}}({\mathbb{R}})$, we obtain
(3.36) ${\big\|\omega_{b}\big\|}_{Y_{\bar{\alpha}}}=$ $\ {\big\|y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}\omega_{b}\big\|}_{L_{x,y}^{2}}+{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\omega_{b}\big\|}_{L_{x,y}^{2}}+{\big\||\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}\omega_{b}\big\|}_{L_{x,y}^{2}}$
$\ \ +{\big\||\partial_{x}|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\omega_{b}\big\|}_{L_{x,y}^{2}}+{\big\||\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}\omega_{b}\big\|}_{L_{x,y}^{2}}$
$\leq$ $\ C{\big\||\partial_{x}|^{\frac{1}{6}+\frac{\bar{\alpha}}{2}}g\big\|}_{L^{2}({\mathbb{R}})}\leq C{\big\||\partial_{x}|^{\frac{13+3\bar{\alpha}}{6}}A\big\|}_{L^{2}({\mathbb{R}})}.$
It follows from (3.35) and (3.36) that $\omega_{b}\in X_{\bar{\alpha}}$ with
${\big\|\omega_{b}\big\|}_{X_{\bar{\alpha}}}\leq C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}.$
Now we are in a position to address the system (3.1), i.e., complete the proof of Theorem 1.1. Let $(\omega_{0},u_{0},v_{0},A_{0})$ be the solution to the linear system
(3.37) $\left\{\begin{aligned} &y\partial_{x}\omega_{0}-\partial_{y}^{2}\omega_{0}=0,\\ &u_{0}=I_{y}[\omega_{0}],\quad v_{0}=-\partial_{x}I_{y}[u_{0}],\\ &\partial_{y}\omega_{0}\big|_{y=0}=\partial_{x}|\partial_{x}|A_{0}(x),\quad I_{M}[\omega_{0}]=A_{0}(x)+F(x).\end{aligned}\right$.
Unless otherwise specified, the parameters $\alpha,\bar{\alpha}$ satisfy $\alpha>0$ and $\bar{\alpha}\geq 0$.
Proof of Theorem 1.1.
We take (3.34) as a solution to (3.37), that is,
(3.38) $\hat{\omega}_{0}(\xi,y)=\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}(i\xi|\xi|\hat{A}_{0})=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}}i\xi|\xi|^{\frac{2}{3}}\hat{A}_{0}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}}i\xi|\xi|^{\frac{2}{3}}\hat{A}_{0}(\xi),\quad\xi<0.\end{aligned}\right$.
Here $\hat{f}(\xi,y)={\mathcal{F}}_{x\to\xi}[f](\xi,y)$. It is straight to verify that
(3.39) $\omega_{0}(\cdot,y)\big|_{y=M}=\frac{Ai(M\partial_{x}^{\frac{1}{3}})}{Ai^{\prime}(0)}\partial_{x}^{\frac{2}{3}}|\partial_{x}|A_{0}.$
Substituting the formula (3.38) into the integral condition $I_{M}[\hat{\omega}_{0}]=\hat{A}_{0}+\hat{F}$, we have
(3.40) $\hat{A}_{0}+\hat{F}$= $\ \int_{0}^{M}\hat{\omega}_{0}(\xi,y)dy$
$=$ $\ \int_{0}^{M}Ai(e^{\frac{{\rm sgn}\ \xi}{6}\pi i}|\xi|^{\frac{1}{3}}y)dy\cdot\frac{1}{Ai^{\prime}(0)e^{\frac{{\rm sgn}\ \xi}{6}\pi i}}i\xi|\xi|^{\frac{2}{3}}\hat{A}_{0}$
$=$ $\ \int_{0}^{e^{\frac{{\rm sgn}\ \xi}{6}\pi i}|\xi|^{\frac{1}{3}}M}Ai(z)dz\cdot\frac{1}{Ai^{\prime}(0)e^{\frac{{\rm sgn}\ \xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}\hat{A}_{0}$
$=$ $\ \frac{p_{M}(\xi)}{Ai^{\prime}(0)e^{\frac{{\rm sgn}\ \xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}\hat{A}_{0}.$
Here
$p_{M}(\xi)=$ $\ \int_{0}^{(i\xi)^{\frac{1}{3}}M}Ai(z)dz.$
Indeed, the integral is well-defined, since $Ai(z)$ is analytic. Let
(3.41) $m(\xi)=$ $1-\frac{p_{M}(\xi)}{Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}.$
Then we get by (3.40) that
(3.42) $\hat{A}_{0}=-\frac{1}{m(\xi)}\hat{F}$.
It follows from (3.38) that
(3.43) $\hat{\omega}_{0}(\xi,y)=-\frac{Ai(e^{\frac{{\rm sgn\ }\xi}{6}\pi i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{6}\pi i}}\frac{i\xi|\xi|^{\frac{2}{3}}}{m(\xi)}\hat{F}(\xi).$
To ensure that the formulas (3.42) and (3.43) are well-defined, we need to prove $m(\xi)\not=0$ for any $\xi\in{\mathbb{R}}$. We define
$p_{0}:=\mathop{sup}\limits_{{\rm Arg\ }z\in\{\pm\frac{\pi}{6}\}}\bigg|\int_{0}^{z}Ai(\zeta)d\zeta\bigg|.$
Noting that $0\leq p_{0}<\infty$, we obtain
$a_{0}:=\bigg(\frac{4p_{0}}{3|Ai^{\prime}(0)|}\bigg)^{-\frac{3}{4}}>0.$
It follows that
(3.44) $m(\xi)\geq\frac{1}{4}\quad\forall\ |\xi|\leq a_{0}.$
Using the following identity
(3.45) $\int_{0}^{\infty}Ai(z)dz=\frac{1}{3},$
the multiplier $m(\xi)$ converges to $\frac{1}{3}$ uniformly on $a_{0},\infty)$ as $M\to\infty$. A direct calculation yields
(3.46) $\bigg|1-\frac{1}{3Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}\bigg|^{-1}\leq C(1+|\xi|)^{-\frac{4}{3}},\quad\forall\ \xi\in{\mathbb{R}}.$
Then there exists a large enough constant $M_{0}>0$ such that for $M\geq M_{0}$, we get by (3.44) and (3.46) that
(3.47) $|m(\xi)|^{-1}=\bigg|1-\frac{p_{M}(\xi)}{Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}\bigg|^{-1}\leq C(1+|\xi|)^{-\frac{4}{3}},\quad\forall\ \xi\in{\mathbb{R}}.$
Thus, $\hat{A}_{0}$ and $\hat{\omega}_{0}$ are well-defined. Moreover, there exists a constant $C>0$ such that
(3.48) ${\big\|m(D)^{-1}f\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big\|f\big\|}_{L^{2}({\mathbb{R}})}.$
Using Proposition 3.2 and (3.48), for any $\bar{\alpha}\geq 0$, we obtain the $X_{\bar{\alpha}}$-estimate of $\omega_{0}$ as follows
${\big\|\omega_{0}\big\|}_{X_{\bar{\alpha}}}\leq C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A_{0}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}F\big\|}_{L^{2}({\mathbb{R}})}\leq C{\big\||\partial_{x}|^{\frac{5}{6}}F\big\|}_{\dot{H}^{\frac{\bar{\alpha}}{2}}({\mathbb{R}})}.$
For the bound of displacement $A_{0}$ and the velocity $u_{0}(x)$, we have
${\big\|u_{0}\big\|}_{L^{\infty}(\Omega_{M})}+{\big\|A_{0}\big\|}_{L^{\infty}({\mathbb{R}})}\leq{\big\|\omega_{0}\big\|}_{L_{x}^{\infty}L_{y}^{1}}\leq{\big\|\omega_{0}\big\|}_{L_{y}^{1}L_{x}^{\infty}}\leq{\big\|\omega_{0}\big\|}_{X_{0}}.$
This completes the proof. ∎
Remark 3.2.
Note that ${\rm Re\ }\bigg(\frac{p_{M}(\xi)}{e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}\bigg)>0$ for $|\xi|^{\frac{1}{3}}M>0$ by numberical results (see figure 2, where $r=|\xi|^{\frac{1}{3}}M$). Then by the continuity of $m(\xi)$, we have
$\frac{1}{|m(\xi)|}\leq C(1+|\xi|)^{-\frac{4}{3}}\text{ for any }\xi\in{\mathbb{R}}\text{ and }M>0,$
where $C>0$ depends on $M^{-1}$.
Figure 2. Analysis on $m(\xi)$.
4. Existence and uniqueness of classical solution
In this section, we establish the existence and uniqueness of classical solution to the nonlinear system (1.13), i.e.,
(4.1) $\left\{\begin{aligned} &y\partial_{x}\omega-\partial_{y}^{2}\omega=-u\partial_{x}\omega-v\partial_{y}\omega,\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\\ &\partial_{y}\omega\big|_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[\omega]=A(x)+F(x).\end{aligned}\right$.
To construct a solution to the system (4.1), we use the iteration method. Let
$(\bar{\omega},\bar{A},\bar{u},\bar{v})=(\omega-\omega_{0},A-A_{0},u-u_{0},v-v_{0}).$
Then the system (4.1) is reduced to the following BVP:
(4.2) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}-\partial_{y}^{2}\bar{\omega}=-(\bar{u}+u_{0})\partial_{x}(\bar{\omega}+\omega_{0})-(\bar{v}+v_{0})\partial_{y}(\bar{\omega}+\omega_{0}),\\ &\bar{u}=I_{y}[\bar{\omega}],\quad\bar{v}=-\partial_{x}I_{y}[\bar{u}],\\ &\partial_{y}\bar{\omega}\big|_{y=0}=\partial_{x}|\partial_{x}|\bar{A},\quad I_{M}[\bar{\omega}]=\bar{A}.\end{aligned}\right$.
We introduce the iterated system as follows
(4.3) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{k+1}-\partial_{y}^{2}\bar{\omega}_{k+1}=-(\bar{u}_{k}+u_{0})\partial_{x}(\bar{\omega}_{k}+\omega_{0})-(\bar{v}_{k}+v_{0})\partial_{y}(\bar{\omega}_{k}+\omega_{0}),\\ &\bar{u}_{k+1}=I_{y}[\bar{\omega}_{k+1}],\quad\bar{v}_{k+1}=-\partial_{x}I_{y}[\bar{u}_{k+1}],\\ &\bar{u}_{k}=I_{y}[\bar{\omega}_{k}],\quad\bar{v}_{k}=-\partial_{x}I_{y}[\bar{u}_{k}],\\ &\partial_{y}\bar{\omega}_{k+1}\big|_{y=0}=\partial_{x}|\partial_{x}|\bar{A}_{k+1},\quad I_{M}[\bar{\omega}_{k+1}]=\bar{A}_{k+1}.\end{aligned}\right$.
First of all, we present the following uniform estimate for the iterated system (4.3).
Proposition 4.1.
Given $\alpha\in(\frac{2}{3},\frac{7}{3}]$, assume that $(\bar{\omega}_{k},\bar{A}_{k})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$. Then there exists a solution $(\bar{\omega}_{k+1},\bar{A}_{k+1})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ to (4.3), which holds that for $\varepsilon\in(0,\frac{1}{6}]$,
(4.4) ${\big\|\bar{\omega}_{k+1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k+1}\big\|}_{X_{\alpha,\infty}}$
$\leq CM^{\frac{5+3\varepsilon}{2}}\Big[\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k}\big\|}_{X_{\alpha,\infty}}\big)^{2}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\Big].$
Here $C=C(\alpha,\varepsilon)>0$ is a constant.
Proof.
We decompose $\bar{\omega}_{k+1}=\bar{\omega}_{e,k+1}+\bar{\omega}_{b,k+1}$, where $\bar{\omega}_{e,k+1}$ and $\bar{\omega}_{b,k+1}$ satisfy
$\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{e,k+1}-\partial_{y}^{2}\bar{\omega}_{e,k+1}=\bar{h}_{k},\\ &\partial_{y}\bar{\omega}_{e,k+1}\big|_{y=0}=0,\quad\bar{\omega}_{e,k+1}\big|_{y=M}=0,\end{aligned}\right$.
and
$\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{b,k+1}-\partial_{y}^{2}\bar{\omega}_{b,k+1}=0,\\ &\partial_{y}\bar{\omega}_{b,k+1}\big|_{y=0}=\partial_{x}|\partial_{x}|\bar{A}_{k+1},\end{aligned}\right$.
together with
(4.5) $I_{M}[\bar{\omega}_{e,k+1}+\bar{\omega}_{b,k+1}]=\bar{A}_{k+1}.$
Here
$\bar{h}_{k}=-(\bar{u}_{k}+u_{0})\partial_{x}(\bar{\omega}_{k}+\omega_{0})-(\bar{v}_{k}+v_{0})\partial_{y}(\bar{\omega}_{k}+\omega_{0}).$
By Proposition 3.1, we obtain the unique solution $\bar{\omega}_{e,k+1}$ with the form
(4.6) $\bar{\omega}_{e,k+1}=$ $\ \pi s_{M}(i\xi)^{-\frac{1}{3}}\bigg\{\mu_{M}(\xi,y)\int_{0}^{y}\mu_{0}(\xi,z)\hat{\bar{h}}_{k}(\xi,z)dz$
$\qquad\qquad\qquad+\mu_{0}(\xi,y)\int_{y}^{M}\mu_{M}(\xi,z)\hat{\bar{h}}_{k}(\xi,z)dz\bigg\}.$
By Proposition 3.2, we can find a solution $\bar{\omega}_{b,k+1}$ with
(4.7) $\hat{\omega}_{b,k+1}(\xi,y)=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),&\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),&xi<0.\end{aligned}\right$.
In view of (4.5), we have
(4.8) $\bar{A}_{k+1}=m(D)^{-1}I_{M}[\bar{\omega}_{e,k+1}].$
Here $m=m(\xi)$ is defined in (3.41).
By Proposition 3.1, Theorem 1.1 and Proposition 2.1, we obtain
(4.9) ${\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}}\leq$ $\ C\big((1+M^{2}){\big\|\bar{h}_{k}\big\|}_{L^{2}(\Omega_{M})}+{\big\||\partial_{x}|^{\frac{\alpha}{2}}\bar{h}_{k}\big\|}_{L^{2}(\Omega_{M})}\big)$
$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}\big({\big\|\bar{\omega}_{0}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|\partial_{x}A_{0}\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}$
$\qquad+{\big\|\partial_{x}A_{k}\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}+{\big\|\partial_{x}F\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}\big)$
$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|A_{k}\big\|}_{X_{\alpha,\infty}}^{2}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\big).$
By Proposition 3.2, Lemma 2.3 and (4.8), we have
(4.10) $\begin{split}{\big\|\bar{A}_{k+1}\big\|}_{X_{\alpha,\infty}}&={\big\||\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}+{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\\ &\leq C{\big\||\partial_{x}|^{\frac{5}{6}}I_{M}[\bar{\omega}_{e,k+1}]\big\|}_{L^{2}({\mathbb{R}})}+C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}I[\bar{\omega}_{e,k+1}]\big\|}_{L^{2}({\mathbb{R}})}\\ &\leq C{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}},\\ {\big\|\bar{\omega}_{b,k+1}\big\|}_{X_{0}\cap X_{\alpha}}&\leq C{\big\||\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}+C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\\ &\leq C{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}},\end{split}$
from which and (4.9), we arrive at
${\big\|\bar{\omega}_{k+1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k+1}\big\|}_{X_{\alpha,\infty}}$
$\leq{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{\omega}_{b,k+1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k+1}\big\|}_{X_{0}\cap X_{\alpha}}$
$\leq C{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}}$
$\leq CM^{\frac{5+3\varepsilon}{2}}\Big[\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|A_{k}\big\|}_{X_{\alpha,\infty}}^{2}\big)+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\Big].$
The proof is completed. ∎
To obtain the uniform estimates of the solution sequence, we use the following basic lemma, which can be proved by the induction.
Lemma 4.1.
Let $\alpha_{i},\beta_{i}\ (i=1,2)$ be positive constants. Assume that $\{a_{k}\}_{k\geq 0},\ \{b_{k}\}_{k\geq 1}$ are two non-negative sequences satisfying
$a_{k+1}\leq\alpha_{1}a_{k}^{2}+\alpha_{2}\quad\mbox{for any }k\geq 0;\quad b_{k+1}\leq\beta_{1}b_{k}+\beta_{2}b_{k}^{2}\quad\mbox{for any }k\geq 1.$
Assume that $\alpha_{2}\leq\frac{1}{4\alpha_{1}}$ and $\beta_{1}<1$. Set $a_{0}=0$ and $b_{1}\leq\frac{\gamma-\beta_{1}}{\beta_{2}}$ for some $\gamma\in(\beta_{1},1)$. Then we have
$0\leq\sup_{k\geq 0}a_{k}\leq 2\alpha_{2},\qquad\sum_{k=1}^{\infty}b_{k}\leq\frac{b_{1}}{1-\gamma}.$
With the above preparations, we complete the proof of Theorem 1.2.
Proof of Theorem 1.2.
The proof consists of three steps.
Step 1. Boundedness of the solution sequence.
Taking $\bar{\omega}_{0}=\bar{A}_{0}=0$, by Proposition 4.1, we obtain a sequence $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$. Set $a_{k}={\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k}\big\|}_{X_{\alpha,\infty}}$. Clearly, we have $a_{0}=0$. By Proposition 4.1, we have
$a_{k+1}\leq CM^{\frac{5+3\varepsilon}{2}}a_{k}^{2}+CM^{\frac{5+3\varepsilon}{2}}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}.$
By Lemma 4.1, for $4C^{2}M^{5+3\varepsilon}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\leq 1$ (that is, ${\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$), we obtain
$\sup_{k\geq 0}\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k}\big\|}_{X_{\alpha,\infty}}\big)=\sup_{k\geq 0}a_{k}\leq 2CM^{\frac{5+3\varepsilon}{2}}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}.$
Step 2. Convergence of the solution sequence.
Set
[…]
By subtracting the equation (4.3) corresponding to $\bar{\omega}_{k+1}$ from the equation (4.3) corresponding to $\bar{\omega}_{k}$, we obtain
[…]
We denote
[…]
In view of Remark 2.1, we have
[…]
It follows from (4.11) and (4.12) that
[…]
By Lemma 4.1, if $CM^{\frac{5+3\varepsilon}{2}}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2}$ and
[…]
which means that ${\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$ for $M\gg 1$, then there holds
[…]
Step 3. Uniqueness of the solution $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$ in ${\mathcal{X}}_{\alpha,M}$.
By Theorem 1.1, we know $(\omega_{0},A_{0})\in{\mathcal{X}}_{\alpha,M}$. Combing (4.6)-(4.7) with the convergence of $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $X_{0}\cap X_{\alpha}\times X_{\alpha,\infty}$, we conclude that $(\bar{\omega},\bar{A})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ and
$\hat{\bar{\omega}}(\xi,y)\big|_{y=M}$ $=\hat{\bar{\omega}}_{b}(\xi,y)\big|_{y=M}$
$=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi<0.\end{aligned}\right$.
This shows that $(\bar{\omega},\bar{A})\in{\mathcal{X}}_{\alpha,M}$. Moreover, it is easy to verify that $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}\setminus\{0\}$ due to $F\not\equiv 0$, which implies that the space ${\mathcal{X}}_{\alpha,M}$ is non-trivial.
Let $(\omega_{i},A_{i})\ (i=1,2)\in{\mathcal{X}}_{\alpha,M}$ be two solutions to the perturbation system (1.13). Set $\tilde{\omega}=\omega_{1}-\omega_{2},\ \tilde{A}=A_{1}-A_{2}$, which solve
[…]
As in the proof of Proposition 4.1, we decompose $\tilde{\omega}=\tilde{\omega}_{e}+\tilde{\omega}_{b}$ with
$\left\{\begin{aligned} &y\partial_{x}\tilde{\omega}_{e}-\partial_{y}^{2}\tilde{\omega}_{e}=\tilde{h},\\ &\partial_{y}\tilde{\omega}_{e}\big|_{y=0}=0,\quad\tilde{\omega}_{e}\big|_{y=M}=0,\end{aligned}\right$.
and
[…]
By Proposition 3.1, we have
(4.13) ${\big\|\tilde{\omega}_{e}\big\|}_{X_{0}\cap X_{\alpha}}\leq$ $\ CM^{2}{\big\|\tilde{h}\big\|}_{L^{2}(\Omega_{M})}+C{\big\|\tilde{h}\big\|}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}}.$
By Proposition 3.2, we get
(4.14) $\hat{\tilde{\omega}}_{b}(\xi,y)=C(\xi)\big(\sqrt{3}Ai((i\xi)^{\frac{1}{3}}y)+Bi((i\xi)^{\frac{1}{3}}y)\big)+\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
In view of $\hat{\tilde{\omega}}_{e}\big|_{y=M}=0$, we have
$\hat{\tilde{\omega}}_{b}\big|_{y=M}=\hat{\tilde{\omega}}\big|_{y=M}=\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)}(i\xi)^{\frac{2}{3}}|\xi|\hat{\tilde{A}}$,
which along with (4.14) gives
$C(\xi)=0.$
Thus, we have
$\hat{\tilde{\omega}}_{b}(\xi,y)=\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
This formula is identical to (3.34). We then get by Proposition 3.2 that
(4.15) ${\big\|\tilde{\omega}_{b}\big\|}_{X_{0}\cap X_{\alpha}}\leq C{\big\||\partial_{x}|^{\frac{5}{6}}\tilde{A}\big\|}_{H^{\frac{4}{3}+\frac{\alpha}{2}}}\leq C{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}.$
Thanks to $I_{M}[\tilde{\omega}_{e}+\tilde{\omega}_{b}]=\tilde{A}$, we have
$\tilde{A}=m(-i\partial_{x})^{-1}I_{M}[\tilde{\omega}_{e}].$
It follows from Lemma 2.3 that
(4.16) ${\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}\leq C{\big\|\tilde{\omega}_{e}\big\|}_{X_{0}\cap X_{\alpha}}.$
By Remark 2.1, we obtain
(4.17) ${\big\|\tilde{h}\big\|}_{L^{2}}\leq\ CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big){\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}$
$\qquad\qquad\quad+CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}\big){\big\|\omega_{2}\big\|}_{X_{0}\cap X_{\alpha}},$
${\big\|\tilde{h}\big\|}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}}\leq CM\big({\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big){\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}$
$\qquad\qquad\quad+CM\big({\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}\big){\big\|\omega_{2}\big\|}_{X_{0}\cap X_{\alpha}}.$
On the other hand, we know that
(4.18) ${\big\|\tilde{A}\big\|}_{L^{\infty}({\mathbb{R}})}\leq{\big\|\tilde{\omega}\big\|}_{L_{x}^{\infty}L_{y}^{1}}\leq{\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}.$
Summing up (4.13), (4.15), (4.16), (4.17) and (4.18), we obtain
${\big\|(\tilde{\omega},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}$
$\leq CM^{\frac{5+3\varepsilon}{2}}{\big\|(\tilde{\omega},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}\big({\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\omega_{2}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big)$
$\leq CM^{\frac{5+3\varepsilon}{2}}{\big\|(\tilde{\omega},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}\big({\big\|(\omega_{1},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|(\omega_{2},A_{2})\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big).$
This implies the uniqueness of solution $(\omega,A)$ in ${\mathcal{X}}_{\alpha,M}$ to (1.13) when ${\big\|(\omega,A)\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\frac{\delta_{0}}{M^{\frac{5+3\varepsilon}{2}}}$ with $\delta_{0}=\delta_{0}(\alpha,\varepsilon)\ll 1$. ∎
Appendix A The physical derivation of triple-deck equation
This appendix provides a brief overview of the triple-deck formalism, based on the physical model of an incompressible boundary layer with a roughness element; for further details, see [26]. The roughness is located at a distance $L$ downstream from the leading edge. Taking $L$ as the reference length, the Reynolds number is defined as $\text{Re}=\frac{U_{\infty}L}{\nu}$, with $U_{\infty}$ and $\nu$ being the velocity of the incoming flow and the kinematic viscosity, respectively. For convenience, we introduce a small parameter $\epsilon=\text{Re}^{-\frac{1}{8}}\ll 1$. The system is described in the Cartesian coordinate $(X,Y)$, with its origin at the leading edge. The roughness height is assumed to be ${\mathcal{O}}(\epsilon^{5})$ or smaller, and its streamwise length scale is ${\mathcal{O}}(\epsilon^{3})$. Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the ${\mathcal{O}}(\epsilon^{3})$ vicinity of the roughness (near $X=1$). The three distinct layers are summarized as follows.
I. Viscous sublayer: Vertical scale ${\mathcal{O}}(\epsilon^{5})$, streamwise scale ${\mathcal{O}}(\epsilon^{3})$.
II. Main layer: Vertical scale ${\mathcal{O}}(\epsilon^{4})$, streamwise scale ${\mathcal{O}}(\epsilon^{3})$.
III. Upper layer: Vertical scale ${\mathcal{O}}(\epsilon^{3})$, streamwise scale ${\mathcal{O}}(\epsilon^{3})$.
These scales arise from matched asymptotic expansions, connecting the triple-deck theory to classical boundary layer theory in the outer flow region. For clarity, the spatial arrangement of layers is illustrated in Figure 3, showing the viscous sublayer I adjacent to the wall and the upper layer III in the potential flow region.
Figure 3. Demonstration of the triple-deck structure (not to scale): I Viscous Sublayer; II Main Layer; III Upper Layer; $\delta$ Thickness of Prandtl layer; $\tilde{U}_{0}$ Streamwise velocity.
It is well-known that the 2D steady incompressible Navier-Stokes equations take the following dimensionless form:
(A.1) $\left\{\begin{aligned} &{\mathbf{u}}_{NS}\cdot\nabla{\mathbf{u}}_{NS}-\epsilon^{8}\Delta{\mathbf{u}}_{NS}+\nabla p_{NS}=0,\\ &\nabla\cdot{\mathbf{u}}_{NS}=0.\end{aligned}\right$.
In the main layer (layer II), we introduce a pair of local coordinates,
$x_{*}=\epsilon^{-3}(X-1)={\mathcal{O}}(1),\quad\tilde{y}=\epsilon^{-4}Y={\mathcal{O}}(1).$
The flow field in this layer can be expressed as a sum of the Blasius base flow ${\mathbf{u}}=(\tilde{U}_{0}(y),0)+{\mathcal{O}}(\epsilon^{4})$ and a mean-flow distortion,
(A.2) $u_{NS}(X,Y;\epsilon)=$ $\ \tilde{U}_{0}(\tilde{y})+\epsilon\tilde{u}(x_{*},\tilde{y})+o(\epsilon),$
$v_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{2}\tilde{v}(x_{*},\tilde{y})+o(\epsilon^{2}),$
$p_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{2}\tilde{p}(x_{*},\tilde{y})+o(\epsilon^{2}).$
Substituting the expansions (A.2) into (A.1) and collecting the leading-order terms, we derive the following system
(A.3) $\left\{\begin{aligned} \tilde{U}_{0}\partial_{x_{*}}\tilde{u}+\tilde{U}_{0}^{\prime}\tilde{v}=&\ 0,\\ \partial_{\tilde{y}}\tilde{p}=&\ 0,\\ \partial_{x_{*}}\tilde{u}+\partial_{\tilde{y}}\tilde{v}=&\ 0.\end{aligned}\right$.
The solutions to this system are
(A.4) $\tilde{u}=\lambda^{-1}A_{*}(x_{*})\tilde{U}_{0}^{\prime}(\tilde{y}),\quad\tilde{v}=-\lambda^{-1}A_{*}^{\prime}(x_{*})\tilde{U}_{0}(\tilde{y}),\quad\tilde{p}=\tilde{p}(x_{*}).$
Here $\lambda:=\tilde{U}_{0}^{\prime}(0)>0$ and $A_{*}$ is an unknown function (typically referred to as a displacement of velocity) satisfying $A_{*}(-\infty)=0$.
As $\tilde{y}\to\infty$, we find that $\tilde{v}$ approaches $\lambda^{-1}A_{*}^{\prime}(x_{*})$, which does not satisfy the far-field attenuation condition, $\lim\limits_{Y\uparrow\infty}v_{NS}(X,Y)=0$. This indicates the emergence of an upper layer (layer III in Figure 3). In this upper layer, we introduce the local coordinate
$\bar{y}=\epsilon^{-3}Y={\mathcal{O}}(1).$
Matching with the upper limit of the main-layer solutions, we can express the flow field as
(A.5) $u_{NS}(X,Y;\epsilon)=$ $1+\epsilon^{2}\bar{u}(x_{*},\bar{y})+o(\epsilon^{2}),$
$v_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{2}\bar{v}(x_{*},\bar{y})+o(\epsilon^{2}),$
$p_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{2}\bar{p}(x_{*},\bar{y})+o(\epsilon^{2}).$
Substituting the expansions (A.5) into (A.1) and retaining the leading-order terms, we derive
[…]
The far-field attenuation condition indicates
(A.7) $\lim_{\sqrt{x_{*}^{2}+\bar{y}^{2}}\to\infty}(\bar{u},\bar{v},\bar{p})=\mathbf{0}.$
Matching with the main layer, we derive
(A.8) $\bar{v}(x_{*},0)=\lim_{\tilde{y}\uparrow\infty}\tilde{v}=-\lambda^{-1}A_{*}^{\prime}(x_{*}),\quad\bar{p}(x_{*},0)=\tilde{p}(x_{*}).$
Since $(\bar{p},\bar{v})$ satisfies the Cauchy-Riemann equations (cf. (A.6)), let us define $f(z)=\bar{p}(x_{*}\bar{y})+i\bar{v}(x_{*},\bar{y})$ with $z=x_{*}+i\bar{y}$. Then, we obtain
(A.9) $f(z)=$ $\ \frac{1}{2\pi i}\int_{{\mathbb{R}}}\frac{f(\zeta)-\overline{f(\zeta)}}{\zeta-z}d\zeta=\ -\frac{1}{\pi\lambda}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-z}d\zeta,\quad\forall\ \text{Im }z>0.$
This implies that
(A.10) $\tilde{p}(x_{*})=\bar{p}(x_{*},0)=\lim_{\text{Im}(z)\to 0^{+}}\text{Re}\ f(z)=-\frac{1}{\pi\lambda}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta.$
Now let us consider the lower boundary of the main-layer solution. As $\tilde{y}\to 0^{+}$, we find that $\tilde{u}$ approaches $A_{*}(x_{*})$, disagree with the no-slip condition. Thus, we need to consider the viscous sublayer (layer I in Figure 3). For convenience, we introduce
$y_{*}=\epsilon^{-5}Y={\mathcal{O}}(1).$
Analysis of the Blasius equation leads to $\tilde{U}_{0}\to\lambda\tilde{y}+{\mathcal{O}}(\tilde{y}^{4})$ as $\tilde{y}\to 0^{+}$. Thus, the main-layer solution behaves like $u_{NS}=\bar{U}_{0}+\epsilon\tilde{u}+o(\epsilon)\to\epsilon(\lambda y_{*}+A_{*})+o(\epsilon)$, which determines the magnitude of the sublayer-layer solution. From the balance of the NS equations, we can expand the solution as,
(A.11) $u_{NS}(X,Y;\epsilon)=$ $\ \epsilon u_{*}(x_{*},y_{*})+o(\epsilon),$
$v_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{3}v_{*}(x_{*},y_{*})+o(\epsilon^{3}),$
$p_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{2}P_{*}(x_{*})+o(\epsilon^{2}).$
Here, we have imposed the fact that, to leading order, the pressure remains uniform across the thin sublayer. Substituting (A.11) into (A.1) and collecting the leading-order terms, we derive the following system:
(A.12) $\left\{\begin{aligned} u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=&\ 0,\\ \partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=&\ 0.\end{aligned}\right$.
The velocity field is subject to the no-slip, non-penetration conditions,
(A.13) $u_{*},v_{*}]\big|_{y_{*}=F(x_{*})}=0.$
In the upstream limit, the velocity approaches the unperturbed Blasius profile,
(A.14) $(u_{*}-\lambda y_{*})\big|_{x_{*}\downarrow-\infty}=0,\quad P_{*}\big|_{x_{*}\downarrow-\infty}=0.$
The upper boundary condition of the sublayer equations should be obtained by matching with the main layer. Now, we consider the overlapping region connecting the sublayer and main layer, where $y_{*}={\mathcal{O}}(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,1)$. Then for any $\beta_{0}>0$, we have
(A.15) $u_{NS}=$ $\ \lambda\tilde{y}+\epsilon\tilde{u}(x_{*},0)+o(\epsilon+\tilde{y}^{4-\beta_{0}})=\lambda\tilde{y}+\epsilon A_{*}(x_{*})+o(\epsilon+\tilde{y}^{4-\beta_{0}})$
$=$ $\ \epsilon\big(\lambda y_{*}+A_{*}(x_{*})\big)+o\big(\epsilon+\epsilon^{(4-\beta_{0})(1-\alpha_{0})}\big).$
To ensure the second part to be negligible, we require $(4-\beta_{0})(1-\alpha_{0})\geq 1$, namely, $\alpha_{0}\in(0,3/4)$ and $\beta_{0}\in(0,4)$. Likewise, we can derive that the leading-order pressure in the sublayer and that in the main layer are equal. Subsequently, we obtain
(A.16) $\big(u_{*}(x_{*},y_{*})-\lambda y_{*}\big)\big|_{y_{*}\uparrow\infty}=A_{*}(x_{*}),\quad\tilde{p}(x_{*})=P_{*}(x_{*}).$
From (A.12)-(A.14) and (A.16), we can derive the classical complete system in viscous sublayer as follows
(A.17) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\quad(x_{*},y_{*})\in\Omega_{\infty,F},\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\quad(x_{*},y_{*})\in\Omega_{\infty,F},\\ &[u_{*},v_{*}]\big|_{y_{*}=F(x_{*})}=0,\quad(u_{*}-\lambda y_{*})\big|_{y_{*}\uparrow\infty}=A_{*}(x_{*}),\quad x\in{\mathbb{R}},\\ &(u_{*}-\lambda y_{*})\big|_{x_{*}\downarrow-\infty}=0,\quad y_{*}\in{\mathbb{R}}_{+},\\ &P_{*}(x_{*})=-\frac{1}{\pi\lambda}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Here $\Omega_{\infty,F}=\big\{(x_{*},y_{*})\big|\ x_{*}\in{\mathbb{R}},\ y_{*}>F(x_{*})\big\}$ and $\lim_{x_{*}\downarrow-\infty}F(x_{*})=0$.
Indeed, the above nonlinear system is usually solved using numerical approach. In light of (A.15), it is permissible to modify the upper boundary of the domain $\Omega_{\infty,F}$ to be a curve $\gamma=\{(x_{*},y_{*})|\ y_{*}=f(x_{*})\in{\mathcal{O}}(\epsilon^{-\alpha_{0}}),\ x_{*}\in{\mathbb{R}}\}$. Consequently, the solution $\phi_{*}=(u_{*},v_{*},P_{*})$ should be denoted by $\phi_{*,\epsilon}$ to include the impact of the upper boundary. It is clear that $\phi_{*,\epsilon}$ also satisfy the system (A.17), except for the upper boundary condition (A.16), which is converted to
(A.18) $(u_{*,\epsilon}(x_{*},y_{*})-\lambda y_{*})\big|_{y_{*}=f(x_{*})}=A_{*}(x_{*}).$
When there is no ambiguity, we still denote $\phi_{*,\epsilon}$ as $\phi_{*}$ for simplicity.
There is a useful operation named the Prandtl transformation for solving the system (A.17), which takes the following form:
(A.19) $y=y_{*}-F(x_{*}),\ v_{*}=V+u_{*}F^{\prime}(x_{*}).$
This transformation merely shifts down the domain by a distance $F(x_{*})$ in the $y$-direction, with the governing equations being unchanged. It is preferable to take $f(x_{*})=\varepsilon^{-\alpha_{0}}+F(x_{*})$, then, the boundary conditions in the transverse direction are changed to
(A.20) $(u_{*},V)\Big|_{y=0}=0,\quad(u_{*}-\lambda y)\Big|_{y=\epsilon^{-\alpha_{0}}}=A_{*}+\lambda F.$
Typically, we take $\lambda=1$ in this paper.
Acknowledgments
M. Dong is partially supported by the NSFC under Grant No. 12588201 and the CAS Project for Young Scientists in Basic Research under Grant No. YSBR-087. C. Wang is partially supported by the NSFC under Grant No. 12471189. Z. Zhang is partially supported by the NSFC under Grant No. 12288101.
Data availability The manuscript has no associated data.
Declarations
Conflict of interest The authors state that there is no conflict of interest.